### v1

# DeepCompile: A Compiler-Driven Approach to Optimizing Distributed Deep Learning Training

Abstract.
The increasing scale of deep learning models has led to the development of various parallelization strategies for distributed training across accelerators. For example, fully sharded approaches like DeepSpeed ZeRO-3 and FSDP partition the parameters of each layer across multiple GPUs and gather them through communication when needed. These methods rely on optimizations such as prefetching, which initiates communication early to overlap it with computation and reduce communication overhead, and unsharding, which retains as many parameters in their unsharded form as possible to reduce communication volume. Although the timing of prefetching should be adjusted in response to dynamic memory usage during execution, these systems lack the flexibility to control it, which limits the benefits of prefetching. Moreover, they cannot anticipate how memory usage will change after prefetching is applied, making it difficult to combine it effectively with other optimizations such as unsharding. We present DeepCompile, which compiles user-defined models into computation graphs and applies a sequence of profiling-guided optimization passes for distributed training. Taking dynamic memory usage into account, these passes flexibly insert, reorder, or remove operations to improve communication–computation overlap, reduce memory pressure, and coordinate multiple optimizations in a unified manner. To evaluate the effectiveness of this design, we implemented a fully sharded approach like ZeRO-3 and FSDP on top of DeepCompile, along with three optimizations: proactive prefetching, selective unsharding, and adaptive offloading. We evaluate DeepCompile on the training of Llama 3 70B and Mixtral 8×7B MoE models. DeepCompile achieves up to 1.28× and 1.54× performance improvements over ZeRO-3 and FSDP baselines, respectively, and up to a 7.01× throughput increase in settings with limited GPU resources, using offloading.
deep learning, distributed training
1. Introduction
The rapid growth of deep learning has led to the emergence of increasingly large models. Modern architectures often contain billions of parameters, resulting in immense computational demands. In this context, efficient parallelization across multiple accelerator devices, such as GPUs, has become essential for achieving feasible training times and manageable costs.
There are several parallelization strategies for training large models: In data parallelism, all model parameters are replicated across multiple GPUs. This approach is natively supported in many deep learning frameworks, including PyTorch (Paszke et al., 2019). Pipeline parallelism (Huang et al., 2019) instead splits the model by layers, assigning each block to a different GPU. Tensor parallelism (Shoeybi et al., 2019) divides parameter tensors within a layer across GPUs, where each device computes a partial result that is later aggregated. In addition to these, the fully sharded approach, which is implemented in DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) (hereafter ZeRO-3) and Fully Sharded Data Parallel (FSDP) (Team, 2023), partitions the parameters of each layer across multiple GPUs and gathers them through communication before each layer’s computation begins.
While these strategies provide scalability, efficiently managing memory and overlapping communication with computation to reduce communication overhead (Wang et al., 2022, 2024) remain open challenges. Existing systems rely on various optimization techniques to address these issues. In the context of the fully sharded approach, such as ZeRO-3 and FSDP, prefetching initiates communication early to gather partitioned parameters and overlap it with computation; unsharding keeps parameters in their full form to reduce communication when memory permits; and offloading, which moves data from GPU memory to host memory, is often used in combination when GPU memory is insufficient.
Although these optimizations are effective in principle, applying them efficiently in practice is not straightforward. Take prefetching as an example. In the fully sharded approach, parameters partitioned across multiple GPUs must be gathered via all-gather communication, which requires a large buffer to hold the gathered parameters. Initiating the all-gather earlier increases opportunities to overlap communication with computation. However, it also extends the buffer’s lifetime and increases memory pressure, since the buffer cannot be released until all computations that use the parameters have completed. Similarly, the overhead of offloading can be mitigated by minimizing data transfers between host and GPU memory and overlapping them with computation. To make these optimizations effective, it is essential to flexibly adjust the timing of all-gather and offloading operations, monitor memory usage trends, and analyze data dependencies across broad scopes such as the forward and backward passes.
Moreover, when multiple optimizations are applied together, their interactions must be carefully coordinated. For instance, unsharding can improve performance when communication is on the critical path, but when combined with prefetching, it must also account for the additional memory pressure introduced by early optimization.
In fully sharded approaches like ZeRO-3 and FSDP, all-gather operations are implemented via runtime hooks inserted into the user-level code (typically in Python). Optimizations such as prefetching and unsharding are supported, but this design lacks the flexibility to adjust the timing of all-gather or offloading operations based on runtime memory usage over broader scopes, such as an entire forward or backward pass. Moreover, there is no clear mechanism to coordinate multiple optimizations in a unified manner.
Other systems have been proposed to automatically combine multiple optimizations for distributed training based on profiling. Examples include Alpa (Mo et al., 2022), FlexFlow (Jia et al., 2018), and Unity (Unger et al., 2022), which aim to determine effective combinations of parallelization strategies such as data, tensor, and pipeline parallelism, as well as to apply kernel-level optimizations like kernel fusion. However, these systems primarily focus on static planning of parallelization strategies and kernel-level optimizations. They do not take into account runtime dynamics such as changes in memory usage during execution. As a result, they lack mechanisms to adjust communication schedules, adapt to memory availability at runtime, or coordinate optimizations such as prefetching, unsharding, and offloading.
To address these limitations, we introduce DeepCompile, a compiler-based system that uses an existing compiler to convert the model into a computation graph and then transforms it for optimization. DeepCompile applies a series of optimization passes, where each pass targets a specific goal—such as improving communication overlap—by inserting, removing, or reordering operations in the graph. These passes make decisions based on profiling such as operator execution time and memory usage trends over broader scopes, including entire forward and backward passes. Since each pass produces a modified graph, subsequent passes can analyze the result and flexibly adapt to the effects of earlier transformations.
To evaluate the effectiveness of this design, we implemented the fully sharded approach, which is similar to ZeRO-3 and FSDP, as a series of optimization passes within DeepCompile. In addition, we implemented the following key optimizations as separate passes.
Proactive prefetching.:
To maximize overlap between communication and computation, this optimization pass initiates all-gather as early as possible, considering how available memory changes as the forward and backward passes progresses.
Selective unsharding.:
This pass keeps as many parameters unsharded as possible to reduce communication overhead caused by all-gather communication, and decides which parameters to unshard based on operator-level memory profiling.
Adaptive offloading.:
DeepCompile offloads optimizer states such as momentum and variance used by Adam (Kingma and Ba, 2014) to CPU memory when GPU memory is insufficient. To reduce data transfer overhead, it offloads only the amount of data that exceeds the memory limit and schedules transfers to overlap with computation.
SimpleFSDP (Zhang et al., 2024) takes a similar approach to DeepCompile by using compilation to optimize the placement of all-gather operations in the fully sharded approach. However, it focuses specifically on prefetching and does not address other optimizations such as unsharding or offloading, which leads to notable differences in design. In SimpleFSDP, communication operations are inserted into the user-level code (Python) before compilation, whereas DeepCompile first compiles the model into a computation graph and then applies transformations. In SimpleFSDP’s design, the compiler must correctly lower these manually inserted communication operations into graph-level operators. While the PyTorch compiler currently supports lowering all-gather operations used by SimpleFSDP, it may fail to handle other optimizations, such as offloading, if they are implemented at the user level. This design limits the extensibility of SimpleFSDP in supporting and combining a broader range of optimizations.
We evaluated DeepCompile on large-scale distributed training tasks using Llama 3 70B and Mixtral 8x7B MoE models. Across all configurations, DeepCompile consistently improved training efficiency compared to baselines, including ZeRO-3, FSDP, and setups using PyTorch compiler optimizations. It achieved up to 1.28× improvement on Llama 3 70B and up to 1.54× on the Mixtral 8x7B MoE model. Additionally, it demonstrated a 7.01× increase in throughput when using fewer GPUs with offloading.
2. Background and Motivating Example
2.1. Distributed Training
The emergence of Transformer architectures (Vaswani et al., 2017), which enabled breakthroughs like BERT (Devlin et al., 2019), has significantly accelerated the growth of modern deep learning models, such as GPT-3 (Brown et al., 2020), Llama-2 (Touvron et al., 2023), Llama-3 (Grattafiori et al., 2024), Phi-3 (Abdin et al., 2024), PaLM (Chowdhery et al., 2022), and OPT (Zhang et al., 2022). This trend has led to a tremendous increase in computational demands. While accelerator devices such as GPUs have become the standard, cutting-edge models now require distributed training across multiple GPUs to be feasible. For example, Llama-3 (Grattafiori et al., 2024) was trained on 16,384 NVIDIA H100 GPUs over 54 days. Even fine-tuning, which adapts pre-trained models to specific downstream tasks, often involves dozens of GPUs running for hours or days.
In distributed training, where multiple GPUs are used in parallel, reducing overhead from both inter-device communication and memory management is critical for efficiently training large-scale models.
2.2. Parallelization Strategies
Various parallelization strategies have been proposed to improve the efficiency of distributed training.
Data parallelism replicates the entire model on each GPU, and each GPU processes a different portion of the training data. A key advantage is that it can be used with any model architecture without requiring code changes, and the communication overhead is relatively low. However, because each GPU must store a full copy of the model, large models may exceed the memory capacity.
Unlike data parallelism, pipeline parallelism partitions the model to fit within GPU memory by grouping layers into coarser-grained stages and placing each stage on a different GPU. The training batch is split into smaller fragments and processed in a pipelined fashion. To fully utilize all GPUs, the compute time of each stage should be approximately balanced. In addition, the choice of partition points affects the communication volume between GPUs. These factors often require manual tuning to achieve optimal performance.
Tensor parallelism is primarily applied to the multi-head attention and MLP layers in Transformer models. It splits specific parameter tensors along designated dimensions across GPUs, performs computations locally, and aggregates results using all-reduce communication. For example, in an MLP block, the first weight matrix is typically split horizontally, and the second one vertically; each GPU computes partial results, and an all-reduce is applied after the second linear transformation to combine them. In addition, although well-tested implementations exist for standard architectures such as GPT (Brown et al., 2020), adapting tensor parallelism to model variants such as multi-query attention (MQA) (Shazeer, 2019), grouped-query attention (GQA) (Narang et al., 2021), or multi-head latent attention (Liu et al., 2024) requires additional engineering effort. In each case, developers must manually identify which parameters can be split, choose appropriate dimensions for partitioning, and ensure correct synchronization across devices.
An alternative to pipeline and tensor parallelism is the fully sharded approach, which partitions the parameters of each layer across multiple GPUs. Frameworks such as Fully Sharded Data Parallel (FSDP) and ZeRO-3 adopt this strategy. During training, each GPU gathers its required parameter shards using all-gather communication before computing each layer. To reduce peak memory usage, the gathered parameters are discarded immediately after the layer’s computation finishes. This approach is widely adopted because it can be applied to arbitrary model architectures without requiring manual restructuring. In this work, we focus on optimizing this fully sharded approach.
2.3. Deep Learning Compiler
Deep learning compilers have rapidly advanced in recent years (Chen et al., 2018; Google, 2023; Ansel et al., 2024). They translate high-level models into computation graphs, which enables graph-level analysis and flexible optimizations such as operator reordering, fusion, and memory reuse.
Major frameworks have been integrating compiler capabilities into their ecosystems. For example, PyTorch now includes its own compiler infrastructure to enable graph-based optimizations (Ansel et al., 2024). Since these compilers can be applied to the vast number of existing model implementations, they are now widely used in practice.
2.4. Motivating Example
The fully sharded approach, as implemented in systems like ZeRO-3 and FSDP, employs runtime optimizations such as prefetching and unsharding. However, these optimizations alone are not sufficient to fully improve training efficiency. Let us take prefetching as an example. Prefetching aims to reduce communication overhead by initiating all-gather operations earlier than the layer where the parameters are actually needed, thereby overlapping communication with computation. Because the fully sharded approach allocates a large buffer to store the gathered parameters, initiating all-gather earlier extends the buffer’s lifetime and increases memory pressure. Therefore, prefetching needs to be scheduled based on memory usage patterns throughout the forward and backward passes.
Figure 1 illustrates a typical pattern of GPU memory usage during a forward and backward pass. Memory usage gradually increases during the forward pass as activations are stored, and decreases during the backward pass as activations are released. At the beginning of the forward pass, a substantial amount of unused memory is available, providing an opportunity to prefetch aggressively and increase computation-communication overlap. In contrast, near the end of the forward pass and throughout the backward pass, memory usage is high, limiting the opportunity for prefetching due to reduced memory availability. As the backward pass progresses and memory is gradually freed, more aggressive prefetching becomes possible again. Existing systems such as ZeRO-3 and FSDP typically allow users to specify a fixed buffer size for prefetching. However, this static configuration cannot adapt to changes in available memory as computation progresses, limiting the potential benefits of prefetching.
Figure 1. Memory usage trends and scheduling opportunities for prefetching and offloading. Profile of several final layers with a sequence length of 4096 and a vocabulary size of 128k. Significant memory spikes are observed in the log-softmax and negative log-likelihood loss layers.
The same issue arises with offloading, which transfers data such as optimizer states (e.g., momentum and variance in Adam optimizer (Kingma and Ba, 2014)) to host memory to reduce GPU memory usage. While offloading can significantly reduce peak GPU memory, the associated data transfers introduce overhead. A naive strategy is to offload states before starting the forward pass, but this leads to significant overhead due to idle GPU time. Instead, data transfers and computation can be overlapped. By initiating transfers at the beginning of the forward pass and synchronizing their completion just before memory usage rises to its limit, it is possible to hide the cost of data movement. In the backward pass, the opposite strategy can be applied: as memory usage decreases, transfers from host memory to GPU can be initiated.
These observations highlight the limitations of existing approaches. To address these limitations, we introduce DeepCompile, a compiler-based approach designed to systematically optimize fully sharded training.
3. System Design
We now describe the design of DeepCompile, a compiler-based system that enables graph-level transformations to support flexible and coordinated optimization of distributed training.
Figure 2 shows the overall workflow. Starting from a user-defined training script written in a framework such as PyTorch, a base compiler (e.g., the PyTorch compiler) lowers the model into an intermediate representation (IR) as a computation graph. DeepCompile takes this graph as input, transforms it by adding communication operations for distributed training, and applies optimizations. The resulting graph is then deployed to a runtime engine running on GPU servers.
Figure 2. Workflow of compilation and optimization with DeepCompile
Note that DeepCompile does not assume that distributed training logic is manually written in the model code. Instead, it targets standard, framework-native implementations, such as those commonly found on the HuggingFace Model Hub [^1] https://huggingface.co/models, and programmatically adds necessary operators including communication and synchronization operations to the graph.
To support dynamic and memory-aware optimizations, DeepCompile organizes its graph transformations as a sequence of optimization passes. Each pass rewrites the computation graph based on a specific strategy, such as inserting communication operators or reordering memory-intensive computations. After applying each pass, DeepCompile executes the modified graph to collect runtime profiling data, including operator execution times, communication overhead, and memory usage trends. This information is then used to guide the next optimization pass. By repeating this process, DeepCompile incrementally refines the graph with awareness of runtime behavior.
In addition, to account for memory usage changes that arise during actual training, DeepCompile periodically runs short training iterations between groups of optimization passes. While DeepCompile focuses primarily on the forward and backward passes, other parts of the training process, such as parameter updates, can significantly affect memory usage. For example, the Adam optimizer typically allocates a large buffer after the first backward pass completes. To capture such changes, DeepCompile runs several training iterations to reflect the updated memory dynamics, then applies another round of optimization passes adapted to the new conditions.
Figure 3 illustrates this two-level loop: an inner loop of optimization and profiling, and an outer loop that periodically runs training to reflect changes in the runtime environment. By alternating between these loops, DeepCompile adapts its transformations to the evolving memory and execution characteristics of the model. This enables coordinated application of optimizations such as prefetching, unsharding, and offloading in a unified and data-driven manner.
Figure 3. Optimization and profiling loop in DeepCompile
While DeepCompile is applicable to a wide range of optimization problems, this paper focuses on one representative use case: the fully sharded approach, as implemented in ZeRO-3 and FSDP. We take advantage of DeepCompile’s graph-level transformation and profiling infrastructure to implement this use case. Specifically, we incorporate a series of optimization passes that target key performance bottlenecks, including proactive prefetching to overlap communication with computation, selective unsharding to reduce redundant data communication, and adaptive offloading to mitigate GPU memory pressure.
In the following sections, we describe the implementation of this fully sharded training strategy using DeepCompile and detail the design and impact of each optimization pass.
4. Optimizations
DeepCompile enables a broad range of optimizations by leveraging a computational graph extracted by the base compiler. In this paper, we first demonstrate how to implement the fully-sharded approach. We then describe additional optimizations to reduce communication overhead and reduce memory requirements.
4.1. Fully-sharded approach
In the fully sharded approach, each parameter tensor is evenly partitioned across all GPUs. Before a layer is computed, each GPU gathers the required parameter shards via an all-gather communication, reconstructing the full parameters locally. Once the layer computation is complete, the gathered parameters are discarded to free memory.
Existing systems such as ZeRO-3 and FSDP insert all-gather and memory release operations around each layer that uses a given parameter. When a model has many layers with small parameter tensors, this results in frequent all-gather operations on small fragments. Since all-gather communication is inefficient for small data sizes (Fig. 5(c)), this leads to significant communication overhead. To mitigate this, both ZeRO-3 and FSDP allow users to manually group multiple layers into a larger block, reducing the number of all-gather operations. However, this increases the lifetime of the buffers holding gathered parameters. Furthermore, none of the layers in the block can begin computation until all of their associated all-gather operations have completed, often resulting in significant idle time. As a result, manually tuning the block size for optimal performance is difficult.
DeepCompile operates by modifying the computation graph to insert all-gather and release operations in a similar manner to existing frameworks. However, unlike systems that rely on layer boundaries or manually defined blocks, DeepCompile uses operator-level dependency analysis to determine the optimal placement of these operations.
As a first step toward enabling the fully sharded approach, DeepCompile schedules all-gather operations just before each parameter’s first use, and release operations immediately after its last use, minimizing the lifetime of each buffer. Figure 4 shows an example of this process. If a parameter is used by multiple operators, DeepCompile analyzes their dependencies to ensure that the buffer is released only after its final use.
Figure 4. Graph modification
Although this scheduling improves memory efficiency, it does not necessarily improve execution speed. DeepCompile addresses this by applying a series of optimization passes. The initial pass focuses on ensuring the model fits within available memory, and subsequent passes improve training efficiency by adjusting the timing of communication, avoiding unnecessary sharding, and overlapping data transfers with computation.
4.2. Proactive prefetching
The initial graph transformation for the fully sharded approach in DeepCompile produces behavior similar to that of ZeRO-3 or FSDP operating in their most fine-grained configuration, where each layer triggers an all-gather before its computation and a release immediately after. However, because DeepCompile expresses this logic at the level of the computation graph, it enables detailed analysis of operator dependencies and memory usage patterns. This graph-based representation provides a foundation for more precise and globally coordinated optimizations, such as prefetching, which go beyond what is possible with heuristic or manually configured approaches.
With the fully sharded approach implemented in the computation graph, DeepCompile can build on top of it by introducing prefetching. Prefetching is a commonly used technique in fully sharded training systems, and both ZeRO-3 and FSDP include built-in support for it. The goal is to improve training efficiency by overlapping communication with computation, reducing the impact of all-gather latency. Figure 5(a) illustrates the timeline of a naive behavior of fully sharded approach, where green boxes represent computation and blue boxes indicate all-gather communication. The number in each blue box corresponds to the respective layer. Since all-gather communication is the primary source of overhead, both ZeRO-3 and FSDP employ prefetching to mitigate its impact. By launching all-gather earlier than the parameters are used, these methods increase opportunities to overlap communication and computation (Fig. 5(b)).
Figure 5. Parameter sharding and prefetching
In existing frameworks, prefetching is implemented in a static and coarse-grained manner. ZeRO-3 issues prefetches based on a predefined buffer size, launching all-gather operations for as many upcoming parameters as can fit in that buffer. Additionally, by grouping multiple layers into a larger block, ZeRO-3 and FSDP can fuse several all-gather operations. This improves communication efficiency, particularly for small data sizes. However, as discussed earlier, it can also delay the computation of earlier layers because the system must wait until all parameters involved in the fused communication are gathered.
Furthermore, ZeRO-3 and FSDP are not designed to account for dynamic changes in memory availability. As shown in Figure 1, memory usage fluctuates throughout the forward and backward passes. More memory tends to be available in the early stages of the forward pass and the later stages of the backward pass. However, existing methods do not effectively take advantage of these patterns. As a result, prefetching opportunities are often underutilized, leading to suboptimal efficiency.
To address this limitation, we propose an optimization pass for proactive prefetching. Figure 6 illustrates the memory usage trend and scheduling behavior with and without proactive prefetching. The goal of this optimization pass is to move all-gather operations as early as possible while ensuring that memory constraints are not violated. Since the memory usage of each operator and the size of the all-gather buffers are known from profiling, it is possible to estimate the total memory footprint if an all-gather is scheduled earlier than usual.
Figure 6. Memory usage and prefetch scheduling decisions. Solid lines indicate memory usage without prefetching; dotted lines show usage with proactive prefetching. All-gather and release operations are labeled accordingly.
In Fig. 6, the dotted line shows the effect of proactive prefetching. In this example, there are three all-gather operations. Based on profiling results, the system determines that prefetching the first two can be safely done at the beginning without exceeding the memory limit. The third all-gather, however, would exceed the memory budget if issued immediately. Therefore, it is scheduled only after the buffer from the first parameter has been released. This scheduling strategy ensures efficient memory usage and allows prefetching to be applied as aggressively as possible within the memory budget.
Algorithm 1 formalizes this procedure. It examines each all-gather operation and determines whether it can be moved earlier in the schedule without exceeding the memory limit, based on the profiled memory usage and buffer size. The algorithm iterates over the operators in the initial schedule $S_{0}$ in reverse order. If the operator $o_{i}$ is an all-gather, it checks whether the operator can be moved earlier without violating the memory constraint.
Table 1. Notations
Symbol Description
$M$ Total memory usage limit
$M_{\text{prefetch}}$ Size limit for the buffer size of a prefetch group
$N$ Number of operators
$S_{0}$ Initial schedule of operators $[o_{1},o_{2},\ldots,o_{N}]$
$OS$ Fragments of optimizer states $[os_{1},os_{2},\ldots]$
$\text{P}_{\text{mem}}(o)$ Profiled memory usage just before operator $o$
$\text{B}_{\text{ag}}(o)$ Buffer size allocated by all-gather operator $o$
$\text{B}_{\text{os}}(os_{i})$ Size of optimizer state fragment $os_{i}$
$Fuse(U)$ Produce operator fusing all-gathers in $U$
$T_{c}(V)$ Communication time for $V$
$\alpha$ Fusion threshold parameter
Algorithm 1 Proactive prefetching
0:  Initial schedule $S_{0}$, Memory limit $M$, Size limit for prefetch group $M_{\text{prefetch}}$.
1:  $S\leftarrow[]$ // Operators in new schedule
2:  $U\leftarrow[]$ // Unscheduled all-gathers
3:  for $i=N$ to $2$ do
4:     if $o_{i}$ is allgather then
5:        $\tilde{m}_{U}\leftarrow\sum_{o\in U\cup\{o_{i}\}}\text{B}_{\text{ag}}(o)$
6:        $\tilde{m}_{i-1}\leftarrow\text{P}_{\text{mem}}(o_{i-1})+\tilde{m}_{U}$
7:        if $\tilde{m}_{i-1}<M$ and $\tilde{m}_{U}<M_{\text{prefetch}}$ then
8:           $U\leftarrow U\cup\{o_{i}\}$ // Add $o_{i}$ to prefetch list
9:        else
10:           $S_{f}\leftarrow\text{Fuse}(U)$, $U\leftarrow[]$
11:           Append $S_{f}$ to $S$ // Schedule fused allgathers
12:        end if
13:     else
14:        Append $o_{i}$ to $S$
15:     end if
16:  end for
17:  $S_{f}\leftarrow\text{Fuse}(U)$
18:  Append $S_{f}$ to $S$ // Schedule all remaining allgathers
19:  return  $S$
If the estimated memory usage remains below the limit, the all-gather operator $o_{i}$ is added to a group $U$ of unscheduled all-gather operators that can be moved earlier. If the memory usage exceeds the limit, the algorithm schedules the all-gather operators in $U$ at the current position, fusing them where appropriate using the Fuse function. This strategy ensures that memory usage stays within the limit $M$.
We also set the limit $M_{\text{prefetch}}$ on the prefetch group, as moving too many allgather operations earlier does not provide any additional benefit. We can set this limit to a significantly larger value than the predefined prefetch buffer size used in existing implementations such as ZeRO-3 because we also ensure the total memory usage is less than the limit $M$. By imposing this limit, we ensure that some memory remains available for selective unsharding, which is discussed in the next section.
When the data size is small, the communication time remains nearly constant, even as the data size increases. Therefore, fusing all-gather operators reduces communication overhead (Fig. 5(c)(d)), especially for small data sizes. The criterion for fusing two all-gather operations is as follows: if $T_{c}(V_{1})+T_{c}(V_{2})>\alpha\cdot T_{c}(V_{1}+V_{2})$, where $V_{1}$ and $V_{2}$ are the data sizes of two operations, fusion is beneficial and the operations are merged. Otherwise, they are left separate. Here, $T_{c}(V)$ is profiled at runtime, and $\alpha$ is a tunable parameter, set to 1.5 in our experiments.
SimpleFSDP (Zhang et al., 2024) takes a similar approach to ours by optimizing prefetching for the fully sharded approach using a compiled computation graph. It performs memory-aware fusion of all-gather operations to reduce communication overhead. However, SimpleFSDP does not consider how memory usage dynamically changes during the forward and backward passes to determine how early each all-gather can be issued. In contrast, our proactive prefetching algorithm uses profiling data to make this decision, allowing communication to be overlapped with computation while staying within memory constraints.
4.3. Selective unsharding
After applying the optimization pass for proactive prefetching, some free memory may remain, since the prefetching size is deliberately limited. Selective unsharding leverages this remaining memory by keeping some parameters unsharded after they are gathered. This approach can significantly reduce communication overhead, although it requires additional memory.
In the fully sharded approach, parameters are distributed across all GPUs, and all-gather operations are required after each parameter update. Selective unsharding, however, allows gathered parameters to remain unsharded after the forward pass, eliminating the need for all-gather in the backward pass.
This technique is especially beneficial when used with gradient accumulation, a common method to increase the effective batch size without increasing per-GPU memory usage. Given a gradient accumulation step of $N$, the model performs $N$ forward and backward passes while accumulating gradients before applying a single parameter update. Since parameters are not updated during this period, they can remain unsharded across multiple forward and backward passes.
We run profiling after applying proactive prefetching to measure the peak memory usage. Then, we select as many parameters to keep unsharded as possible, ensuring that the total buffer size does not exceed the memory limit. The selection considers both the buffer size of each all-gather operator $o$, denoted as $\text{B}(o)$, and its communication time $T_{c}(\text{B}(o))$. Parameters are prioritized based on the ratio $\frac{T_{c}(\text{B}(o))}{\text{B}(o)}$, where a higher value indicates a greater reduction in communication time relative to memory cost. Following the trend shown in Fig. 5(d), smaller parameters are generally selected first.
4.4. Adaptive offloading
Although optimizer states such as momentum and variance used in Adam optimizer (Kingma and Ba, 2014) are only needed for parameter updates and are not required during the forward and backward passes, they occupy a significant amount of memory throughout training. Existing frameworks, such as DeepSpeed, provide functionality to offload these states to CPU memory. In fully sharded approaches like ZeRO-3, both model parameters and optimizer states are partitioned across GPUs. This design reduces the memory footprint on each device and enables parallel data transfers between GPUs and CPU memory. As a result, optimizer state offloading is often used together with fully sharded training.
Algorithm 2 Adaptive offloading (forward)
0:  Initial schedule $S_{0}$, Memory limit $M$, optimizer state fragments $OS$
  $S\leftarrow[]$ // List of operators in output schedule
  $M^{-}\leftarrow 0$ // Offloaded size
  $M_{\text{peak}}\leftarrow max_{o_{i}\in S_{0}}\text{P}_{\text{mem}}(o_{i})$
  $M_{opt}\leftarrow\sum_{os_{i}\in OS}\text{B}_{\text{os}}(os_{i})$
  $OS_{\text{offload}}\leftarrow[]$
  for all $os_{i}\in OS$ do
     if $M_{\text{peak}}+M_{\text{opt}}-\sum_{os_{i}\in OS_{\text{offload}}}\text{B}_{\text{os}}(os_{i})>M$ then
        $OS_{\text{offload}}\leftarrow OS_{\text{offload}}\cup os_{i}$
        Append offload operator for $os_{i}$ to $S$
     end if
  end for
  for all $o_{i}\in S_{0}$ do
     while $\text{P}_{\text{mem}}(o_{i})+M_{\text{opt}}-M^{-}>M$ do
        Pop $os_{i}$ from $OS_{\text{offload}}$
        Append operator to synchronize the copy of $os_{i}$ to $S$ and free memory of $os_{i}$
        $M^{-}\leftarrow M^{-}+\text{B}_{\text{os}}(os_{i})$
     end while
     Append $o_{i}$ to $S$
  end for
  return  $S$
However, offloading and reloading optimizer states can typically significant overhead. To address this challenge, we propose an approach called adaptive offloading, which improves upon existing offloading mechanisms in two ways. First, by monitoring actual memory usage during training, it minimizes the amount of data that needs to be offloaded. Only the portion of optimizer states that would exceed the memory limit is transferred to the host. Second, it takes advantage of the characteristic memory usage patterns in training. Since memory usage tends to increase during the forward pass and decrease during the backward pass, adaptive offloading schedules data transfers to overlap with computation. This overlap reduces the performance impact of moving data between GPU and CPU memory.
Algorithm 2 presents the pseudo-code for the forward pass with adaptive offloading of optimizer states. The optimizer states are divided into a large enough number of fragments. At the start of the forward pass, the algorithm schedules operators to initiate asynchronous offloading of fragments that exceed the memory limit. Since the copy operation is asynchronous, computation can run in parallel. The memory must only be freed after the copy operation completes. Therefore, the peak memory usage before each operator is checked based on profiling when scheduling. If the memory usage exceeds the limit, the algorithm adds operators to synchronize the copy and free memory for the fragments.
We also iterate over the operators in the backward pass schedule, checking the available memory. If sufficient memory is available from that operator to the end, we schedule a reload operation to transfer the optimizer states back to the GPU. This approach enables overlapping the backward computation with the data transfer from host memory to GPU memory while adhering to memory constraints. Since memory usage often increases during forward passes and decreases during backward passes, this strategy allows for efficient overlap of computation and offloading/reloading operations.
5. Evaluation
We conducted experiments to evaluate the performance of DeepCompile across multiple dimensions, including computational efficiency, memory utilization, and correctness.
5.1. Experimental settings
The experiments were performed on two or four servers, each equipped with eight NVIDIA H100 GPUs (80GB memory per GPU) and 1TB of system memory. The GPUs within each node were interconnected via NVLink, and each server included eight network interfaces (HCAs) connected via InfiniBand. The nodes were interconnected using InfiniBand with an observed bandwidth of 340GB/s for both allgather and reduce-scatter operations. Each server was equipped with two Intel(R) Xeon(R) Platinum 8462Y+ processors.
In our experiments, we used the Llama-3 70B and Mixtral 8x7B models (47 billion parameters)(Jiang et al., 2024) as our target workloads. These models are representative examples of dense and Mixture-of-Experts (MoE) architectures(Shazeer et al., 2017; Lepikhin et al., 2021; Fedus et al., 2022), respectively. Both were run using bfloat16 with mixed precision (Micikevicius et al., 2017), while maintaining FP32 copies of parameters, gradients, and Adam optimizer states. We also employed activation checkpointing, integrated into the model implementation at the Transformer layer level, where each layer is recomputed as a block.
We used PyTorch v2.6.0 and models from the HuggingFace Model Hub without any built-in parallelization support. The experiments were conducted with Python 3.10 and CUDA 12.4. DeepSpeed v1.6.4 and PyTorch FSDP were used as the baseline frameworks, and DeepCompile was built on top of DeepSpeed. The optimization passes and profilers in DeepCompile were implemented in Python (approximately 2.6K lines), while the custom operators injected into the computational graph were implemented in C++ (approximately 900 lines).
While some distributed training strategies support partial sharding—for example, sharding only optimizer states, or both gradients and optimizer states while keeping parameters unsharded—these configurations are not sufficient for our target models. The total parameter sizes of Llama-3 70B and Mixtral 8x7B exceed the memory capacity of a single GPU (140 GB and 94 GB, respectively), requiring all parameters, gradients, and optimizer states to be fully sharded across devices. For this reason, we use the fully sharded configurations of both DeepSpeed ZeRO-3 and PyTorch FSDP in our evaluation.
Other large-scale training frameworks, such as Megatron-LM, are not included in our comparison. Although Megatron-LM has been widely adopted for large model training, it requires a specialized implementation where parallelism strategies are tightly coupled with model code. In contrast, we focus on model-agnostic approaches that can be applied to standard implementations. Similarly, frameworks that automate parallelization, such as Alpa (Mo et al., 2022), FlexFlow (Jia et al., 2018), and Unity (Unger et al., 2022), are not included, as they remain experimental and do not currently support the models used in this evaluation.
5.2. Efficiency
Figure 7 presents the throughputs observed in our experiments. We used Llama-3-70B with 32 GPUs and Mixtral-8x7B with 16 GPUs. We evaluated DeepSpeed ZeRO-3 both without and with the PyTorch compiler enabled, labeled as ZeRO3 and ZeRO3 (C), respectively. Similarly, we include results for PyTorch FSDP, labeled as FSDP and FSDP (C) depending on whether the PyTorch compiler is enabled. For DeepCompile, we tested three configurations: enabling only proactive prefetching, enabling only selective unsharding, and enabling both in that order. These configurations are labeled as DeepCompile (P), DeepCompile (S), and DeepCompile (P+S), respectively, in the charts. Since the models fit within the available GPU memory at these scales, adaptive offloading was not applied in these experiments.
We conducted this experiment using different batch sizes and sequence lengths. In addition, we evaluated training performance with gradient accumulation.
Figure 7. Throughputs resulting from Llama-3 70B and Mixtral 8x7B models
DeepCompile consistently outperformed all baseline methods, with the most significant gains observed when both proactive prefetching and selective unsharding were enabled (DeepCompile (P+S)). Under the setting with a gradient accumulation step of 1, the smallest improvement over ZeRO-3 was 1.11×, observed in the configuration with sequence length 512 and batch size 1. In this setting, computation is significantly faster relative to communication, leaving little opportunity for our optimizations to hide communication overhead through overlap. As a result, the benefit of prefetching and unsharding becomes limited. For all other conditions, DeepCompile achieved approximately 1.14× to 1.18× speedup over ZeRO-3, and 1.16× to 1.26× over FSDP in the Llama-3-70B experiments. For Mixtral-8x7B, the speedups over ZeRO-3 ranged from 1.07× to 1.35×, while those over FSDP ranged from 1.06× to 1.14×. Mixtral-8x7B is an MoE model, where only a subset of parameters is activated for each input token. Due to its relatively small computation load compared to total parameter size, it suffers more from communication bottlenecks. As a result, similar to Llama-3-70B, the improvements were slightly larger when the batch size and sequence length were increased, enabling better overlap between communication and computation.
Figure 7(a)(iii) and (b)(iii) show throughput results when varying the gradient accumulation steps from 1 to 16. As discussed in Section 4.3, this setting significantly benefits from selective unsharding. In the Llama-3-70B experiment (Fig. 7(a)(iii)), DeepCompile achieves increasingly higher throughput as the accumulation step increases, with up to 1.28× improvement over ZeRO-3 at step 16. In Mixtral-8x7B (Fig. 7(b)(iii)), the improvement is even more pronounced due to the model’s heavier communication load. DeepCompile achieves up to 1.54× higher throughput compared to ZeRO-3 at step 16.
In contrast, FSDP fails to run with accumulation steps greater than 1, as it does not support accumulating gradients while keeping them partitioned. Instead, it gathers full gradients after each backward pass, which causes the total memory usage to exceed GPU capacity, leading to memory allocation failures.
5.3. Memory Utilization
The optimization passes described above allow DeepCompile to utilize GPU memory more effectively, contributing to increased throughput. We compare the memory utilization of DeepCompile against other baselines. Figure 8 presents the peak GPU memory usage for different models across various sequence lengths.
Figure 8. Memory utilization from Llama-3 70B and Mixtral 8x7B models
The memory footprints of ZeRO-3 and FSDP are approximately 40GB for Llama-3-70B and 30GB for Mixtral-8x7B. When selective unsharding is enabled, DeepCompile actively utilizes available memory to keep more parameters unsharded. We configure a safety margin by first reserving approximately 7GB for the CUDA driver, NCCL buffers, and other runtime components, and then applying an additional 10% margin to the remaining memory. As a result, DeepCompile (S) and DeepCompile (P+S) consistently use around 65GB in all cases. This demonstrates that the selective unsharding mechanism dynamically adapts the amount of unsharded parameters based on the available memory at runtime.
5.4. Adaptive Offloading
The Llama-3-70B model requires 32 or more GPUs, each with 80GB of memory, to fit entirely into GPU memory under our current settings. However, by offloading optimizer states, we are able to fit the model onto just 16 GPUs. To evaluate the benefits of adaptive offloading of optimizer states, as described in Section 4.4, we compared the throughput achieved using DeepSpeed’s offloading feature with that of our adaptive offloading approach.
Figure 9 shows the iteration times for each method. ZeRO3 (Offload optimizer) and ZeRO3 (Offload optimizer) + Compile represent the results of using DeepSpeed ZeRO-3 with and without the PyTorch compiler, respectively. Notably, DeepSpeed not only offloads optimizer states to host memory but also performs parameter updates on the CPU. While this approach saves more memory than DeepCompile’s offloading, it significantly slows down parameter updates. Since this behavior makes it difficult to compare DeepSpeed’s offloading directly with our method—which keeps parameter updates on the GPU—we also included results for DeepCompile (Offload optimizer all + sync), which offloads all optimizer states at the beginning of the forward pass and reloads them synchronously at the end. Our adaptive offloading method is shown as DeepCompile (Offload selective + async), which minimizes the amount of data offloaded and overlaps offloading and reloading asynchronously with computation.
The experimental results in Figure 9 indicate that adaptive offloading (DeepCompile (Offload selective + async)) achieves up to 7.0× and 6.9× higher throughput than ZeRO3 (Offload optimizer) and ZeRO3 (Offload optimizer) + Compile, respectively, when the batch size is 1 and the sequence length is 1024. It also demonstrates a 2.6× improvement over DeepCompile (Sync).
Figure 9. Throughput results for adaptive offloading
5.5. Compilation time
As shown in Fig.3, DeepCompile applies a sequence of optimization passes. In our experiments, we first executed the pass that inserts allgather and release operations. After a short warmup period (five training iterations), we applied additional passes including proactive prefetching and selective unsharding.
These passes involve graph analysis and profiling, resulting in several minutes of overhead for our target models. Table 2 reports the time required for each configuration, measured with batch size 1 and sequence length 512.
Table 2. Time for compilation
Proactive prefetching Selective unsharding Both
Llama-3 70B 254.8 s 248.6 s 266.9 s
Mixtral 8x7B 437.0 s 416.5 s 407.8 s
The variation in compile time across optimization configurations for the same model is less than 10%, indicating that the additional cost of enabling proactive prefetching or selective unsharding is relatively minor compared to the base analysis and scheduling pass. On the other hand, compile times vary more significantly across models, as seen in the longer durations for Mixtral-8x7B compared to Llama-3-70B. This difference stems from model-specific factors such as graph size and layer composition.
It is important to note that this compilation overhead occurs only once before training begins. During actual training, the compiled graph is reused without incurring any additional overhead. Given that model training typically consists of hundreds to thousands of iterations, the one-time compilation cost is negligible in practice.
Furthermore, the results of the optimization passes can be cached and reused across training runs as long as key settings such as batch size and sequence length remain unchanged. This reuse is particularly valuable when sweeping hyperparameters such as learning rates or dropout ratios, enabling users to avoid repeating the initial compilation and further reducing end-to-end overhead.
5.6. Correctness
To verify the correctness of the optimization passes in DeepCompile, we compared the resulting loss values with those of DeepSpeed ZeRO-3. We initialized Llama-3 70B model with random weights but the same random seed in both settings. We used the AG News corpus (Zhang et al., 2016) as training examples, with the sequence length and micro batch size set to 512 and 1, respectively. The learning rate was set to 1.5e-5. Figure 10 shows the training losses for both settings. Although some operators are non-deterministic and introduce subtle differences, the loss curves were closely aligned.
Figure 10. Verification of loss values
6. Related Work
6.1. Distributed Training Framework
As discussed in Section 2, various distributed training strategies have been proposed to scale deep learning models beyond the capacity of a single GPU. For instance, tensor parallelism, as introduced in Megatron-LM (Shoeybi et al., 2019), splits large matrix operations across devices, while pipeline parallelism, as implemented in GPipe (Huang et al., 2019), assigns consecutive layers to different GPUs. While effective, both approaches require the user to manually modify the model to specify how tensors or layers are partitioned. This makes it challenging to adapt these strategies to diverse or rapidly evolving model architectures.
In contrast, our work focuses on the fully sharded approach, as implemented in ZeRO (Rajbhandari et al., 2020) and FSDP (Team, 2023), which requires no model code changes and is applicable to a wide range of architectures. DeepCompile builds on this fully sharded approach and introduces compiler-level optimizations to improve its efficiency.
Note that DeepCompile is not inherently limited to the fully sharded approach. Its design is compatible with other parallelization strategies. For example, RaNNC (Tanaka et al., 2021) automatically partitions the computation graph to enable pipeline parallelism, and Alpa (Mo et al., 2022) selects distributed execution rules based on the partitioned state of tensors. Similarly, DeepCompile could support automated parallelization by replacing computation operators in the graph and integrating this with other optimization passes within a unified framework.
6.2. Optimizing Parallelization Strategies
To improve training efficiency for large-scale models, various strategies such as data, tensor, and pipeline parallelism have been studied. Recent work has focused on automatically discovering optimal combinations of these strategies, including Alpa (Mo et al., 2022), FlexFlow (Jia et al., 2018), Unity (Unger et al., 2022), and Galvatron (Miao et al., 2022). Dynamic switching techniques have also been proposed, such as HotSpa (Ge et al., 2024), which aims to adaptively reconfigure parallelization strategies during training.
However, these approaches generally assume that the model is already implemented in a way that supports the desired forms of parallelism. In practice, tensor and pipeline parallelism often require model-specific partitioning logic and integration with low-level APIs. Supporting multiple strategies and enabling dynamic switching can involve considerable engineering effort, particularly for complex or evolving model architectures. As a result, despite their promise, such automated parallelization systems have seen limited adoption in current practice.
6.3. Computation-Communication Overlapping
As discussed earlier, widely used frameworks such as ZeRO-3 and FSDP incorporate prefetching techniques to overlap communication with computation and reduce latency from all-gather operations. Prefetching has become a standard optimization for improving efficiency in fully sharded training.
More recently, techniques for overlapping communication at a finer granularity, such as intra-layer parallelism, have also been proposed (Wang et al., 2022, 2024). These approaches aim to eliminate or hide communication costs by restructuring computation within layers. In contrast, our work focuses on profiling-guided, graph-level optimization of communication overlap.
Like DeepCompile, SimpleFSDP (Zhang et al., 2024) targets compiler-level optimization of the fully sharded approach. It performs memory-aware fusion of all-gather operations to reduce communication overhead, but does not take into account dynamic changes in memory usage when deciding how early each all-gather can be issued without exceeding memory constraints. Moreover, SimpleFSDP focuses solely on scheduling communication for FSDP. In contrast, DeepCompile is designed to support a broader range of optimizations and enables techniques such as selective unsharding and adaptive offloading.
7. Conclusion
We introduced DeepCompile, a compiler-based framework for optimizing distributed training of large-scale deep learning models. DeepCompile starts from a standard single-GPU computation graph and incrementally inserts distributed operators such as all-gather and release, enabling graph-level analysis and holistic optimization of distributed training. Unlike existing frameworks like ZeRO-3 and FSDP, which rely on heuristics and fixed module boundaries, DeepCompile leverages runtime profiling to flexibly reorder and coordinate operations based on actual memory usage and data dependencies.
Built on this foundation, DeepCompile implements a fully sharded training approach along with three key optimizations: proactive prefetching to improve communication–computation overlap, selective unsharding to reduce redundant communication, and adaptive offloading to dynamically manage memory under tight resource constraints. These optimization passes are applied in sequence, informed by profiling feedback, allowing DeepCompile to coordinate them effectively in a unified compilation framework.
Our experiments demonstrate that DeepCompile consistently outperforms ZeRO-3 and FSDP baselines across a range of configurations. It achieves up to 1.28× and 1.54× higher throughput for Llama-3 70B and Mixtral 8×7B models, respectively. Moreover, in scenarios with limited GPU memory, DeepCompile’s adaptive offloading achieves up to a 7.01× throughput improvement by automatically overlapping communication and reducing memory pressure.
DeepCompile represents a step forward in compiler-driven distributed training. Its graph transformation capabilities and runtime-aware optimization passes offer a scalable and automated path toward efficient parallelization. In future work, we aim to extend DeepCompile to support broader classes of optimizations, such as fully automated parallelization planning and enhanced memory scheduling.

### v2

@@
 Abstract.
-The increasing scale of deep learning models has led to the development of various parallelization strategies for distributed training across accelerators.
-For example, fully sharded approaches like DeepSpeed ZeRO-3 and FSDP partition the parameters of each layer across multiple GPUs and gather them through communication when needed.
-These methods rely on optimizations such as prefetching, which initiates communication early to overlap it with computation and reduce communication overhead, and unsharding, which retains as many parameters in their unsharded form as possible to reduce communication volume.
-Although the timing of prefetching should be adjusted in response to dynamic memory usage during execution, these systems lack the flexibility to control it, which limits the benefits of prefetching.
-Moreover, they cannot anticipate how memory usage will change after prefetching is applied, making it difficult to combine it effectively with other optimizations such as unsharding.
-We present DeepCompile, which compiles user-defined models into computation graphs and applies a sequence of profiling-guided optimization passes for distributed training.
-Taking dynamic memory usage into account, these passes flexibly insert, reorder, or remove operations to improve communication–computation overlap, reduce memory pressure, and coordinate multiple optimizations in a unified manner.
-To evaluate the effectiveness of this design, we implemented a fully sharded approach like ZeRO-3 and FSDP on top of DeepCompile, along with three optimizations: proactive prefetching, selective unsharding, and adaptive offloading.
-We evaluate DeepCompile on the training of Llama 3 70B and Mixtral 8×7B MoE models.
-DeepCompile achieves up to 1.28× and 1.54× performance improvements over ZeRO-3 and FSDP baselines, respectively, and up to a 7.01× throughput increase in settings with limited GPU resources, using offloading.
+The rapid growth of deep learning models has increased the demand for efficient distributed training strategies.
+Fully sharded approaches like ZeRO-3 and FSDP partition model parameters across GPUs and apply optimizations such as prefetching and unsharding to reduce communication overhead.
+However, these systems lack fine-grained control over memory and communication scheduling, making it difficult to balance computation–communication overlap with memory requirements.
+Coordinating multiple optimizations such as prefetching and unsharding is also difficult, since their effects on memory usage can influence each other.
+To tackle these challenges, we propose DeepCompile, a compiler-based optimization framework for distributed training.
+DeepCompile transforms user-defined models into computation graphs and applies a series of profiling-guided optimization passes, each modifying the graph based on profiling information such as execution time and memory usage.
+This design allows each pass to flexibly insert, reorder, or remove operations such as all-gather and memory allocation, improving communication–computation overlap and reducing memory pressure.
+Each pass can access updated profiling feedback from earlier passes, enabling coordinated optimizations.
+We further enhance DeepCompile by three additional optimizations: proactive prefetching, selective unsharding, and adaptive offloading.
+Our evaluation shows that DeepCompile achieves up to 1.28× and 1.54× speedups over ZeRO-3 and FSDP baselines, respectively, and up to a 7.01× throughput increase in settings with limited GPU resources using offloading.
 deep learning, distributed training
@@
 The rapid growth of deep learning has led to the emergence of increasingly large models.
-Modern architectures often contain billions of parameters, resulting in immense computational demands.
-In this context, efficient parallelization across multiple accelerator devices, such as GPUs, has become essential for achieving feasible training times and manageable costs.
-There are several parallelization strategies for training large models: In data parallelism, all model parameters are replicated across multiple GPUs.
+Modern architectures often contain tens or even hundreds of billions of parameters (Brown et al., 2020; Zhang et al., 2022; Abdin et al., 2024; Grattafiori et al., 2024; Chowdhery et al., 2022), resulting in immense computational demands.
+While deep learning compilers (Chen et al., 2018; Google, 2023; Ansel et al., 2024) have been highly successful in improving training efficiency on a single accelerator device, such as a GPU, efficient parallelization across multiple devices has become essential to fit large models into memory.
+There are several parallelization strategies for training large models.
+In data parallelism, all model parameters are replicated across multiple GPUs.
 This approach is natively supported in many deep learning frameworks, including PyTorch (Paszke et al., 2019).
@@
 Tensor parallelism (Shoeybi et al., 2019) divides parameter tensors within a layer across GPUs, where each device computes a partial result that is later aggregated.
-In addition to these, the fully sharded approach, which is implemented in DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) (hereafter ZeRO-3) and Fully Sharded Data Parallel (FSDP) (Team, 2023), partitions the parameters of each layer across multiple GPUs and gathers them through communication before each layer’s computation begins.
-While these strategies provide scalability, efficiently managing memory and overlapping communication with computation to reduce communication overhead (Wang et al., 2022, 2024) remain open challenges.
-Existing systems rely on various optimization techniques to address these issues.
-In the context of the fully sharded approach, such as ZeRO-3 and FSDP, prefetching initiates communication early to gather partitioned parameters and overlap it with computation; unsharding keeps parameters in their full form to reduce communication when memory permits; and offloading, which moves data from GPU memory to host memory, is often used in combination when GPU memory is insufficient.
-Although these optimizations are effective in principle, applying them efficiently in practice is not straightforward.
-Take prefetching as an example.
-In the fully sharded approach, parameters partitioned across multiple GPUs must be gathered via all-gather communication, which requires a large buffer to hold the gathered parameters.
-Initiating the all-gather earlier increases opportunities to overlap communication with computation.
-However, it also extends the buffer’s lifetime and increases memory pressure, since the buffer cannot be released until all computations that use the parameters have completed.
-Similarly, the overhead of offloading can be mitigated by minimizing data transfers between host and GPU memory and overlapping them with computation.
-To make these optimizations effective, it is essential to flexibly adjust the timing of all-gather and offloading operations, monitor memory usage trends, and analyze data dependencies across broad scopes such as the forward and backward passes.
-Moreover, when multiple optimizations are applied together, their interactions must be carefully coordinated.
-For instance, unsharding can improve performance when communication is on the critical path, but when combined with prefetching, it must also account for the additional memory pressure introduced by early optimization.
-In fully sharded approaches like ZeRO-3 and FSDP, all-gather operations are implemented via runtime hooks inserted into the user-level code (typically in Python).
-Optimizations such as prefetching and unsharding are supported, but this design lacks the flexibility to adjust the timing of all-gather or offloading operations based on runtime memory usage over broader scopes, such as an entire forward or backward pass.
-Moreover, there is no clear mechanism to coordinate multiple optimizations in a unified manner.
-Other systems have been proposed to automatically combine multiple optimizations for distributed training based on profiling.
-Examples include Alpa (Mo et al., 2022), FlexFlow (Jia et al., 2018), and Unity (Unger et al., 2022), which aim to determine effective combinations of parallelization strategies such as data, tensor, and pipeline parallelism, as well as to apply kernel-level optimizations like kernel fusion.
-However, these systems primarily focus on static planning of parallelization strategies and kernel-level optimizations.
-They do not take into account runtime dynamics such as changes in memory usage during execution.
-As a result, they lack mechanisms to adjust communication schedules, adapt to memory availability at runtime, or coordinate optimizations such as prefetching, unsharding, and offloading.
-To address these limitations, we introduce DeepCompile, a compiler-based system that uses an existing compiler to convert the model into a computation graph and then transforms it for optimization.
-DeepCompile applies a series of optimization passes, where each pass targets a specific goal—such as improving communication overlap—by inserting, removing, or reordering operations in the graph.
-These passes make decisions based on profiling such as operator execution time and memory usage trends over broader scopes, including entire forward and backward passes.
-Since each pass produces a modified graph, subsequent passes can analyze the result and flexibly adapt to the effects of earlier transformations.
-To evaluate the effectiveness of this design, we implemented the fully sharded approach, which is similar to ZeRO-3 and FSDP, as a series of optimization passes within DeepCompile.
-In addition, we implemented the following key optimizations as separate passes.
-Proactive prefetching.:
-To maximize overlap between communication and computation, this optimization pass initiates all-gather as early as possible, considering how available memory changes as the forward and backward passes progresses.
-Selective unsharding.:
-This pass keeps as many parameters unsharded as possible to reduce communication overhead caused by all-gather communication, and decides which parameters to unshard based on operator-level memory profiling.
-Adaptive offloading.:
-DeepCompile offloads optimizer states such as momentum and variance used by Adam (Kingma and Ba, 2014) to CPU memory when GPU memory is insufficient.
-To reduce data transfer overhead, it offloads only the amount of data that exceeds the memory limit and schedules transfers to overlap with computation.
-SimpleFSDP (Zhang et al., 2024) takes a similar approach to DeepCompile by using compilation to optimize the placement of all-gather operations in the fully sharded approach.
-However, it focuses specifically on prefetching and does not address other optimizations such as unsharding or offloading, which leads to notable differences in design.
-In SimpleFSDP, communication operations are inserted into the user-level code (Python) before compilation, whereas DeepCompile first compiles the model into a computation graph and then applies transformations.
-In SimpleFSDP’s design, the compiler must correctly lower these manually inserted communication operations into graph-level operators.
-While the PyTorch compiler currently supports lowering all-gather operations used by SimpleFSDP, it may fail to handle other optimizations, such as offloading, if they are implemented at the user level.
-This design limits the extensibility of SimpleFSDP in supporting and combining a broader range of optimizations.
+In addition to these, the fully sharded approach, which is implemented in DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) (hereafter ZeRO-3) and Fully Sharded Data Parallel (FSDP) (Team, 2023), partitions the parameters of each layer across multiple GPUs and gathers them through all-gather communication before each layer’s computation begins.
+Although existing parallelization strategies have enabled large-scale model training, two key challenges remain unsolved:
+Challenge 1: Fine-grained control over communication and memory scheduling.
+Existing systems often lack the ability to precisely control when to initiate communication or allocate and release memory buffers.
+As a result, it is difficult to optimize the overlap between communication and computation, or to efficiently manage memory usage that fluctuates during execution.
+Challenge 2: Coordinating multiple optimizations.
+As existing deep learning compilers apply a sequence of independent optimization passes, such as kernel fusions and operator reordering, distributed training also requires independent optimizations such as communication scheduling and memory planning.
+However, there is currently no unified mechanism to combine these optimizations and handle their interactions effectively.
+To illustrate Challenge 1, consider prefetching in the fully sharded approach.
+Initiating all-gather communication earlier can improve overlap with computation, but prefetching too much data can exhaust available memory, as the buffers cannot be released until all dependent computations are completed.
+A representative case of Challenge 2 involves combining prefetching with unsharding, which retains as many parameters in their unsharded form as possible to reduce communication overhead.
+To combine these two optimizations effectively, the system must accurately estimate how much memory remains available after prefetching is applied.
+Existing systems partially address these challenges.
+Runtime-based approaches such as ZeRO-3 and FSDP support optimizations like prefetching and unsharding, but lack fine-grained control over the timing of communication and memory scheduling in response to runtime memory changes.
+Compiler-based systems such as Alpa (Mo et al., 2022), Unity (Unger et al., 2022), nnScalar (Lin et al., 2024) automatically combine different parallelization strategies, but do not adapt to dynamic memory behavior during execution.
+To bridge this gap, we introduce DeepCompile, a compiler-based optimization framework for distributed training.
+DeepCompile uses an existing compiler to convert a model into a computation graph and applies a series of transformations to optimize multi-GPU execution.
+These transformations improve communication overlap and reduce memory pressure based on profiling information that captures execution time and memory usage trends across the forward and backward passes.
+This approach addresses the two key challenges discussed earlier:
+Graph transformations enabling fine-grained memory-aware scheduling.
+By directly transforming the computation graph, DeepCompile can flexibly inject, reorder, and remove operations such as all-gather.
+This fine-grained control enables precise placement of communication and memory operations, which is difficult to express in user-level code.
+Furthermore, by capturing operator-level memory usage patterns across forward and backward passes, DeepCompile identifies optimal points to initiate communication and allocate or release memory based on runtime memory availability.
+Composable and unified optimization passes with profiling feedback.
+DeepCompile organizes optimizations into modular optimization passes, each transforming the computation graph based on profiling data such as execution time and memory usage.
+While each pass can operate sequentially, later passes can access updated profiling data from the transformed graph to observe and respond to the effects of earlier passes.
+This design enables multiple distributed training optimizations to be composed in a globally coordinated manner.
+To evaluate the effectiveness of this design, we implemented the fully sharded approach in DeepCompile as an optimization pass, similar in functionality to ZeRO-3 (Rajbhandari et al., 2020) and FSDP (Team, 2023)4.1).
+This pass demonstrates that the fully sharded paradigm can be captured within the optimization pass framework, and serves as a foundation for building additional compiler-based optimizations.
+Building on this foundation, we design and incorporate three additional optimization passes into DeepCompile.
+Each pass targets a distinct aspect of communication or memory optimization in fully sharded training.
+  • Proactive prefetching (§4.2) initiates all-gather operations earlier to improve overlap between communication and computation, while taking into account dynamic memory availability during execution.
+  • Selective unsharding (§4.3) avoids sharding for selected parameters based on operator-level memory profiling, reducing all-gather operations for sharded parameters and adapting to memory usage trends.
+  • Adaptive offloading (§4.4) transfers optimizer states, such as the momentum and variance buffers used by the Adam optimizer (Kingma and Ba, 2014), to CPU memory when GPU memory is insufficient.
+These transfers are scheduled to overlap with computation to hide latency.
+All of these passes are applied after the optimization pass enabling the baseline fully sharded functionality described in §4.1.
+Moreover, as we show in later experiments, combining proactive prefetching and selective unsharding yields significant performance improvements over applying either pass in isolation.
+SimpleFSDP (Zhang et al., 2024) takes a similar approach to DeepCompile by using compilation to improve the placement of all-gather operations in the fully sharded approach.
+However, it is specialized for FSDP and focuses solely on prefetching, along with fused all-gather communications.
+As a result, it does not support composing multiple modular optimization passes.
+Coordination across passes using profiling feedback—for example, adapting later optimizations to the effects of earlier ones—is not considered in its design.
 We evaluated DeepCompile on large-scale distributed training tasks using Llama 3 70B and Mixtral 8x7B MoE models.
@@
 It achieved up to 1.28× improvement on Llama 3 70B and up to 1.54× on the Mixtral 8x7B MoE model.
-Additionally, it demonstrated a 7.01× increase in throughput when using fewer GPUs with offloading.
+Additionally, it delivered a 7.01× increase in throughput in low-GPU settings where, even with parameter partitioning, the model does not fit in GPU memory without offloading.
 2. Background and Motivating Example
-2.1. Distributed Training
-The emergence of Transformer architectures (Vaswani et al., 2017), which enabled breakthroughs like BERT (Devlin et al., 2019), has significantly accelerated the growth of modern deep learning models, such as GPT-3 (Brown et al., 2020), Llama-2 (Touvron et al., 2023), Llama-3 (Grattafiori et al., 2024), Phi-3 (Abdin et al., 2024), PaLM (Chowdhery et al., 2022), and OPT (Zhang et al., 2022).
-This trend has led to a tremendous increase in computational demands.
-While accelerator devices such as GPUs have become the standard, cutting-edge models now require distributed training across multiple GPUs to be feasible.
-For example, Llama-3 (Grattafiori et al., 2024) was trained on 16,384 NVIDIA H100 GPUs over 54 days.
-Even fine-tuning, which adapts pre-trained models to specific downstream tasks, often involves dozens of GPUs running for hours or days.
-In distributed training, where multiple GPUs are used in parallel, reducing overhead from both inter-device communication and memory management is critical for efficiently training large-scale models.
+The rapid scaling of Transformer-based architectures (Vaswani et al., 2017) has driven a dramatic increase in model size and computational demands.
+While accelerators like GPUs are standard, training state-of-the-art models now requires large-scale distributed training.
+For instance, Llama-3 (Grattafiori et al., 2024) was trained on 16,384 NVIDIA H100 GPUs over 54 days.
+Even fine-tuning tasks frequently span dozens of GPUs.
+As such, reducing overhead from inter-device communication and memory usage has become essential for efficient training.
 2.1. Parallelization Strategies
-Various parallelization strategies have been proposed to improve the efficiency of distributed training.
-Data parallelism replicates the entire model on each GPU, and each GPU processes a different portion of the training data.
-A key advantage is that it can be used with any model architecture without requiring code changes, and the communication overhead is relatively low.
-However, because each GPU must store a full copy of the model, large models may exceed the memory capacity.
-Unlike data parallelism, pipeline parallelism partitions the model to fit within GPU memory by grouping layers into coarser-grained stages and placing each stage on a different GPU.
-The training batch is split into smaller fragments and processed in a pipelined fashion.
-To fully utilize all GPUs, the compute time of each stage should be approximately balanced.
-In addition, the choice of partition points affects the communication volume between GPUs.
-These factors often require manual tuning to achieve optimal performance.
-Tensor parallelism is primarily applied to the multi-head attention and MLP layers in Transformer models.
-It splits specific parameter tensors along designated dimensions across GPUs, performs computations locally, and aggregates results using all-reduce communication.
-For example, in an MLP block, the first weight matrix is typically split horizontally, and the second one vertically; each GPU computes partial results, and an all-reduce is applied after the second linear transformation to combine them.
-In addition, although well-tested implementations exist for standard architectures such as GPT (Brown et al., 2020), adapting tensor parallelism to model variants such as multi-query attention (MQA) (Shazeer, 2019), grouped-query attention (GQA) (Narang et al., 2021), or multi-head latent attention (Liu et al., 2024) requires additional engineering effort.
-In each case, developers must manually identify which parameters can be split, choose appropriate dimensions for partitioning, and ensure correct synchronization across devices.
+To address the computational and memory demands of large-scale models, several parallelization strategies have been developed.
+Data parallelism replicates the entire model on each GPU and splits the input data.
+While easy to adopt, it cannot scale to very large models that exceed single-GPU memory.
+Pipeline parallelism splits the model by layers into sequential stages across GPUs.
+It reduces memory usage but requires careful balancing of stage workloads and often manual tuning.
+Tensor parallelism partitions individual weight matrices across devices and aggregates partial results using all-reduce.
+Though effective for standard architectures with well-supported implementations (e.g., GPT (Brown et al., 2020)), it often requires non-trivial engineering to adapt to variants such as multi-query attention (Shazeer, 2019) or grouped-query attention (Ainslie et al., 2023).
 An alternative to pipeline and tensor parallelism is the fully sharded approach, which partitions the parameters of each layer across multiple GPUs.
-Frameworks such as Fully Sharded Data Parallel (FSDP) and ZeRO-3 adopt this strategy.
-During training, each GPU gathers its required parameter shards using all-gather communication before computing each layer.
+Frameworks such as DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) and Fully Sharded Data Parallel (FSDP) (Team, 2023) adopt this strategy.
+In this approach, each GPU gathers its required parameter shards using all-gather communication before computing each layer.
 To reduce peak memory usage, the gathered parameters are discarded immediately after the layer’s computation finishes.
@@
 Deep learning compilers have rapidly advanced in recent years (Chen et al., 2018; Google, 2023; Ansel et al., 2024).
-They translate high-level models into computation graphs, which enables graph-level analysis and flexible optimizations such as operator reordering, fusion, and memory reuse.
+They translate user-defined models into computation graphs, which enables graph-level analysis and flexible optimizations such as operator reordering, fusion, and memory reuse.
 Major frameworks have been integrating compiler capabilities into their ecosystems.
@@
 Therefore, prefetching needs to be scheduled based on memory usage patterns throughout the forward and backward passes.
-Figure 1 illustrates a typical pattern of GPU memory usage during a forward and backward pass.
-Memory usage gradually increases during the forward pass as activations are stored, and decreases during the backward pass as activations are released.
+Figure 1 illustrates a typical pattern of GPU memory usage during forward and backward passes.
+Memory usage gradually increases during the forward pass as activations are stored for the following backward pass, and decreases during the backward pass as activations are released.
 At the beginning of the forward pass, a substantial amount of unused memory is available, providing an opportunity to prefetch aggressively and increase computation-communication overlap.
@@
 Profile of several final layers with a sequence length of 4096 and a vocabulary size of 128k.
-Significant memory spikes are observed in the log-softmax and negative log-likelihood loss layers.
-The same issue arises with offloading, which transfers data such as optimizer states (e.g., momentum and variance in Adam optimizer (Kingma and Ba, 2014)) to host memory to reduce GPU memory usage.
-While offloading can significantly reduce peak GPU memory, the associated data transfers introduce overhead.
-A naive strategy is to offload states before starting the forward pass, but this leads to significant overhead due to idle GPU time.
+Significant memory spikes are observed in the log-softmax and negative log-likelihood loss layers in the backward pass.
+The same issue arises with offloading, which transfers data such as optimizer states (e.g., momentum and variance in the Adam optimizer (Kingma and Ba, 2014)) to host memory to reduce GPU memory usage during forward and backward passes.
+This is possible because optimizer states are only needed after the backward pass, when updating parameters.
+Though offloading can significantly reduce peak GPU memory usage, the associated data transfers introduce overhead.
+A naive strategy is to offload optimizer states before the forward pass and reload them after the backward pass finishes, since they are not used during computation.
+However, this leads to idle GPU time during data transfers.
 Instead, data transfers and computation can be overlapped.
-By initiating transfers at the beginning of the forward pass and synchronizing their completion just before memory usage rises to its limit, it is possible to hide the cost of data movement.
+By initiating transfers at the beginning of the forward pass and synchronizing their completion just before memory usage reaches its peak, it is possible to hide the cost of data movement.
 In the backward pass, the opposite strategy can be applied: as memory usage decreases, transfers from host memory to GPU can be initiated.
-These observations highlight the limitations of existing approaches.
-To address these limitations, we introduce DeepCompile, a compiler-based approach designed to systematically optimize fully sharded training.
+To achieve this, the system must accurately track how memory usage changes during execution and place offload, reload, and synchronization operations at appropriate points.
+These observations motivate the design of DeepCompile, a compiler-based approach for systematically optimizing fully sharded training.
 3. System Design
@@
 Figure 2 shows the overall workflow.
-Starting from a user-defined training script written in a framework such as PyTorch, a base compiler (e.g., the PyTorch compiler) lowers the model into an intermediate representation (IR) as a computation graph.
+Starting from a user-defined training script and a model implementation written in a framework such as PyTorch, a base compiler (e.g., the PyTorch compiler) lowers the model into an intermediate representation (IR) as a computation graph.
 DeepCompile takes this graph as input, transforms it by adding communication operations for distributed training, and applies optimizations.
 The resulting graph is then deployed to a runtime engine running on GPU servers.
-Figure 2. Workflow of compilation and optimization with DeepCompile
+Figure 2. Workflow of DeepCompile. DeepCompile transforms the IR generated by a base compiler, injecting distributed training optimizations before execution on GPUs.
 Note that DeepCompile does not assume that distributed training logic is manually written in the model code.
 Instead, it targets standard, framework-native implementations, such as those commonly found on the HuggingFace Model Hub [^1] https://huggingface.co/models, and programmatically adds necessary operators including communication and synchronization operations to the graph.
-To support dynamic and memory-aware optimizations, DeepCompile organizes its graph transformations as a sequence of optimization passes.
+To support dynamic and memory-aware optimizations, DeepCompile organizes its graph transformations as a sequence of optimization passes, as is common in existing deep learning compilers.
 Each pass rewrites the computation graph based on a specific strategy, such as inserting communication operators or reordering memory-intensive computations.
@@
 To capture such changes, DeepCompile runs several training iterations to reflect the updated memory dynamics, then applies another round of optimization passes adapted to the new conditions.
-Figure 3 illustrates this two-level loop: an inner loop of optimization and profiling, and an outer loop that periodically runs training to reflect changes in the runtime environment.
+Figure 3 illustrates this two-level loop: an inner loop of optimization and profiling for each optimization pass, and an outer loop that periodically runs training to reflect changes in the runtime environment.
 By alternating between these loops, DeepCompile adapts its transformations to the evolving memory and execution characteristics of the model.
-This enables coordinated application of optimizations such as prefetching, unsharding, and offloading in a unified and data-driven manner.
-Figure 3. Optimization and profiling loop in DeepCompile
+This enables the coordination of multiple optimizations such as prefetching, unsharding, and offloading in a unified and profile-guided manner.
+Figure 3.
+Profiling-guided optimization loop in DeepCompile.
+DeepCompile applies optimizations in iterative phases, with periodic training to reflect memory dynamics and coordinate passes such as prefetching and offloading.
 While DeepCompile is applicable to a wide range of optimization problems, this paper focuses on one representative use case: the fully sharded approach, as implemented in ZeRO-3 and FSDP.
@@
 DeepCompile enables a broad range of optimizations by leveraging a computational graph extracted by the base compiler.
-In this paper, we first demonstrate how to implement the fully-sharded approach.
-We then describe additional optimizations to reduce communication overhead and reduce memory requirements.
+In this paper, we begin by describing how the fully sharded approach is expressed within our framework, and then introduce additional optimizations that reduce communication overhead and memory requirements.
 4.1. Fully-sharded approach
@@
 Once the layer computation is complete, the gathered parameters are discarded to free memory.
-Existing systems such as ZeRO-3 and FSDP insert all-gather and memory release operations around each layer that uses a given parameter.
-When a model has many layers with small parameter tensors, this results in frequent all-gather operations on small fragments.
-Since all-gather communication is inefficient for small data sizes (Fig.
-5(c)), this leads to significant communication overhead.
-To mitigate this, both ZeRO-3 and FSDP allow users to manually group multiple layers into a larger block, reducing the number of all-gather operations.
-However, this increases the lifetime of the buffers holding gathered parameters.
-Furthermore, none of the layers in the block can begin computation until all of their associated all-gather operations have completed, often resulting in significant idle time.
-As a result, manually tuning the block size for optimal performance is difficult.
-DeepCompile operates by modifying the computation graph to insert all-gather and release operations in a similar manner to existing frameworks.
-However, unlike systems that rely on layer boundaries or manually defined blocks, DeepCompile uses operator-level dependency analysis to determine the optimal placement of these operations.
+Figure 4. Inserting all-gather and release operations for the fully sharded approach. This minimizes buffer lifetime by aligning communication and memory operations with actual usage.
+In existing implementations of the fully sharded approach, such as ZeRO-3 and FSDP, communication and memory operations are typically triggered by hooks inserted at layer boundaries in the Python module structure.
+These hooks invoke all-gather operations and manage memory allocation and release in response to the forward and backward pass.
+In contrast, DeepCompile operates by modifying the computation graph to insert all-gather and release operations, similar to existing frameworks.
+However, unlike systems that rely on layer boundaries or manually defined blocks, DeepCompile can flexibly inject or reorder such communication and memory operations based on the dependencies among operators.
 As a first step toward enabling the fully sharded approach, DeepCompile schedules all-gather operations just before each parameter’s first use, and release operations immediately after its last use, minimizing the lifetime of each buffer.
@@
 If a parameter is used by multiple operators, DeepCompile analyzes their dependencies to ensure that the buffer is released only after its final use.
-Figure 4. Graph modification
 Although this scheduling improves memory efficiency, it does not necessarily improve execution speed.
-DeepCompile addresses this by applying a series of optimization passes.
-The initial pass focuses on ensuring the model fits within available memory, and subsequent passes improve training efficiency by adjusting the timing of communication, avoiding unnecessary sharding, and overlapping data transfers with computation.
+To address this, DeepCompile applies a series of optimization passes.
+The initial pass ensures that the model fits within available memory, while subsequent passes improve training efficiency by adjusting the timing of communication, avoiding unnecessary sharding, and overlapping data transfers with computation.
 4.2. Proactive prefetching
-The initial graph transformation for the fully sharded approach in DeepCompile produces behavior similar to that of ZeRO-3 or FSDP operating in their most fine-grained configuration, where each layer triggers an all-gather before its computation and a release immediately after.
-However, because DeepCompile expresses this logic at the level of the computation graph, it enables detailed analysis of operator dependencies and memory usage patterns.
-This graph-based representation provides a foundation for more precise and globally coordinated optimizations, such as prefetching, which go beyond what is possible with heuristic or manually configured approaches.
-With the fully sharded approach implemented in the computation graph, DeepCompile can build on top of it by introducing prefetching.
-Prefetching is a commonly used technique in fully sharded training systems, and both ZeRO-3 and FSDP include built-in support for it.
-The goal is to improve training efficiency by overlapping communication with computation, reducing the impact of all-gather latency.
-Figure 5(a) illustrates the timeline of a naive behavior of fully sharded approach, where green boxes represent computation and blue boxes indicate all-gather communication.
-The number in each blue box corresponds to the respective layer.
-Since all-gather communication is the primary source of overhead, both ZeRO-3 and FSDP employ prefetching to mitigate its impact.
-By launching all-gather earlier than the parameters are used, these methods increase opportunities to overlap communication and computation (Fig.
-5(b)).
-Figure 5. Parameter sharding and prefetching
-In existing frameworks, prefetching is implemented in a static and coarse-grained manner.
-ZeRO-3 issues prefetches based on a predefined buffer size, launching all-gather operations for as many upcoming parameters as can fit in that buffer.
-Additionally, by grouping multiple layers into a larger block, ZeRO-3 and FSDP can fuse several all-gather operations.
-This improves communication efficiency, particularly for small data sizes.
-However, as discussed earlier, it can also delay the computation of earlier layers because the system must wait until all parameters involved in the fused communication are gathered.
-Furthermore, ZeRO-3 and FSDP are not designed to account for dynamic changes in memory availability.
-As shown in Figure 1, memory usage fluctuates throughout the forward and backward passes.
-More memory tends to be available in the early stages of the forward pass and the later stages of the backward pass.
-However, existing methods do not effectively take advantage of these patterns.
-As a result, prefetching opportunities are often underutilized, leading to suboptimal efficiency.
-To address this limitation, we propose an optimization pass for proactive prefetching.
-Figure 6 illustrates the memory usage trend and scheduling behavior with and without proactive prefetching.
-The goal of this optimization pass is to move all-gather operations as early as possible while ensuring that memory constraints are not violated.
-Since the memory usage of each operator and the size of the all-gather buffers are known from profiling, it is possible to estimate the total memory footprint if an all-gather is scheduled earlier than usual.
+The initial graph transformation for the fully sharded approach in DeepCompile produces behavior similar to that of ZeRO-3 and FSDP, where each layer triggers an all-gather before its computation and releases the gathered parameters immediately afterward.
+Figure 5(a) illustrates this baseline behavior during a forward pass.
+Computation (green boxes) and all-gather communication (blue boxes) are executed sequentially per layer, resulting in no overlap.
+Figure 5.
+Impact of prefetching on memory usage and communication–computation overlap.
+(a) Without prefetching, all-gather operations are issued just before computation, resulting in no overlap.
+(b) Prefetching enables overlap, but fixed buffer sizes may cause out-of-memory errors.
+A dynamic strategy adapts to memory availability and avoids such failures.
+The bottom track in each timeline shows memory usage over time.
+Activation memory (green), used for intermediate computation, accumulates across layers and is retained until the backward pass.
+In contrast, memory used for all-gather communication (blue) increases when the operation begins and is released once the corresponding computation completes.
+As a result, more memory is typically available earlier in the forward pass and later in the backward pass.
+To reduce communication overhead, systems such as ZeRO-3 and FSDP support prefetching, which launches all-gather operations earlier than strictly required to enable communication–computation overlap.
+Figure 5(b) demonstrates how different prefetching strategies impact memory usage and execution behavior.
+With a fixed prefetching buffer size of 2, safe overlap is achieved without exceeding memory limits.
+However, a larger buffer size of 4 causes an out-of-memory error due to increased memory pressure from overlapping communication and accumulating activation memory.
+To prevent such failures, ZeRO-3 and FSDP employ a static buffer size and limits prefetching to the parameters that fit within that bound.
+While safe, this conservative strategy fails to adapt to dynamic changes in memory availability during training.
+As illustrated in Figure 1, memory usage fluctuates significantly throughout execution.
+Fine-grained control is required to determine exactly when to initiate each all-gather operation.
+The bottom timeline in Figure 5(b) exemplifies such dynamic scheduling, which is not feasible in existing systems like ZeRO-3 or FSDP.
+To address this limitation, we introduce an optimization pass for proactive prefetching, which schedules all-gather operations as early as possible while ensuring that memory constraints are not violated.
+By profiling memory usage and all-gather buffer sizes after applying the optimization pass for the fully sharded approach, DeepCompile can estimate the total memory footprint of earlier communication.
 Figure 6.
@@
 All-gather and release operations are labeled accordingly.
-In Fig.
-6, the dotted line shows the effect of proactive prefetching.
-In this example, there are three all-gather operations.
-Based on profiling results, the system determines that prefetching the first two can be safely done at the beginning without exceeding the memory limit.
-The third all-gather, however, would exceed the memory budget if issued immediately.
-Therefore, it is scheduled only after the buffer from the first parameter has been released.
-This scheduling strategy ensures efficient memory usage and allows prefetching to be applied as aggressively as possible within the memory budget.
-Algorithm 1 formalizes this procedure.
-It examines each all-gather operation and determines whether it can be moved earlier in the schedule without exceeding the memory limit, based on the profiled memory usage and buffer size.
-The algorithm iterates over the operators in the initial schedule $S_{0}$ in reverse order.
-If the operator $o_{i}$ is an all-gather, it checks whether the operator can be moved earlier without violating the memory constraint.
+Figure 6 illustrates how this scheduling works.
+Solid lines indicate memory usage without prefetching, while dotted lines show usage with proactive prefetching.
+In this example, three all-gather operations are scheduled.
+Based on the memory budget, the first two can be safely prefetched at the beginning.
+The third, however, would exceed the memory limit if issued immediately, so it is delayed until after the buffer for the first parameter has been released.
+This approach maximizes prefetching opportunities while maintaining safe memory usage throughout execution.
 Table 1. Notations
 Symbol Description
+$S_{0}$ Initial schedule of operators $o_{1},o_{2},\ldots,o_{N}]$
 $M$ Total memory usage limit
 $M_{\text{prefetch}}$ Size limit for the buffer size of a prefetch group
-$N$ Number of operators
-$S_{0}$ Initial schedule of operators $[o_{1},o_{2},\ldots,o_{N}]$
-$OS$ Fragments of optimizer states $[os_{1},os_{2},\ldots]$
-$\text{P}_{\text{mem}}(o)$ Profiled memory usage just before operator $o$
+NN Number of operators
+OSOS Fragments of optimizer states $os_{1},os_{2},\ldots]$
+$\text{P}_{\text{mem}}(o)$ Profiled memory usage before operator $o$
 $\text{B}_{\text{ag}}(o)$ Buffer size allocated by all-gather operator $o$
@@
 0: Initial schedule $S_{0}$, Memory limit $M$, Size limit for prefetch group $M_{\text{prefetch}}$.
-1:  $S\leftarrow[]$ // Operators in new schedule
-2:  $U\leftarrow[]$ // Unscheduled all-gathers
-3:  for $i=N$ to $2$ do
+[…]
+2:U[$\leftarrow[] // U$nscheduled all-gathers
+3:for $i=N$ to 22 do
 4:  if $o_{i}$ is allgather then
@@
 16:end for
-17:  $S_{f}\leftarrow\text{Fuse}(U)$
+[…]
 18: Append $S_{f}$ to $S$ // Schedule all remaining allgathers
 19:return $S$
+Algorithm 1 formalizes this procedure.
+It iterates over the operators in the initial schedule $S_{0}$ in reverse order.
+The schedule $S_{0}$ is produced by the preceding optimization pass that enables the fully sharded approach.
+For each operator $o_{i}$, if it is an all-gather operation, the algorithm checks whether it can be moved earlier in the schedule without violating the memory constraint.
+This decision is based on the profiled memory usage and the required buffer size.
 If the estimated memory usage remains below the limit, the all-gather operator $o_{i}$ is added to a group $U$ of unscheduled all-gather operators that can be moved earlier.
-If the memory usage exceeds the limit, the algorithm schedules the all-gather operators in $U$ at the current position, fusing them where appropriate using the Fuse function.
-This strategy ensures that memory usage stays within the limit $M$.
-We also set the limit $M_{\text{prefetch}}$ on the prefetch group, as moving too many allgather operations earlier does not provide any additional benefit.
-We can set this limit to a significantly larger value than the predefined prefetch buffer size used in existing implementations such as ZeRO-3 because we also ensure the total memory usage is less than the limit $M$.
+If the usage exceeds the limit, the algorithm schedules the operators in $U$ at the current position, applying fusion based on the Fuse function.
+This strategy maintains memory usage within the limit $M$.
+We also set the limit $M_{\text{prefetch}}$ on the prefetch group, as moving too many all-gather operations earlier does not provide any additional benefit.
 By imposing this limit, we ensure that some memory remains available for selective unsharding, which is discussed in the next section.
-When the data size is small, the communication time remains nearly constant, even as the data size increases.
-Therefore, fusing all-gather operators reduces communication overhead (Fig.
-5(c)(d)), especially for small data sizes.
-The criterion for fusing two all-gather operations is as follows: if $T_{c}(V_{1})+T_{c}(V_{2})>\alpha\cdot T_{c}(V_{1}+V_{2})$, where $V_{1}$ and $V_{2}$ are the data sizes of two operations, fusion is beneficial and the operations are merged.
-Otherwise, they are left separate.
-Here, $T_{c}(V)$ is profiled at runtime, and $\alpha$ is a tunable parameter, set to 1.5 in our experiments.
-SimpleFSDP (Zhang et al., 2024) takes a similar approach to ours by optimizing prefetching for the fully sharded approach using a compiled computation graph.
+The Fuse function decides whether to fuse a set of all-gather operators and returns a sequence of fused or unfused operations.
+For small message sizes, communication time increases slowly and remains nearly flat.
+As a result, combining multiple small all-gather operations into a single larger one can significantly reduce communication overhead.
+The $\text{Fuse}(U)$ operation fuses two all-gather calls if the following condition holds: if $T_{c}(V_{1})+T_{c}(V_{2})>\alpha\cdot T_{c}(V_{1}+V_{2})$, where $V_{1}$ and $V_{2}$ are the data sizes of two operations.
+Here, $T_{c}(V)$ is profiled communication time for size $V$, and $\alpha$ is a tunable parameter.
+SimpleFSDP (Zhang et al., 2024) takes a similar approach to ours to optimize the fully sharded approach using a compiled computation graph.
 It performs memory-aware fusion of all-gather operations to reduce communication overhead.
@@
 4.3. Selective unsharding
-After applying the optimization pass for proactive prefetching, some free memory may remain, since the prefetching size is deliberately limited.
-Selective unsharding leverages this remaining memory by keeping some parameters unsharded after they are gathered.
-This approach can significantly reduce communication overhead, although it requires additional memory.
-In the fully sharded approach, parameters are distributed across all GPUs, and all-gather operations are required after each parameter update.
-Selective unsharding, however, allows gathered parameters to remain unsharded after the forward pass, eliminating the need for all-gather in the backward pass.
-This technique is especially beneficial when used with gradient accumulation, a common method to increase the effective batch size without increasing per-GPU memory usage.
-Given a gradient accumulation step of $N$, the model performs $N$ forward and backward passes while accumulating gradients before applying a single parameter update.
-Since parameters are not updated during this period, they can remain unsharded across multiple forward and backward passes.
+After applying the optimization pass for proactive prefetching, some free memory may remain.
+Selective unsharding leverages this remaining memory by keeping certain parameters in their unsharded form after they are gathered and delaying their release until the parameter update step, which occurs after the backward pass.
+By avoiding repeated all-gather operations for these parameters, this approach reduces communication overhead.
+This technique is especially beneficial when used with gradient accumulation, a common method to increase the effective batch size without increasing GPU memory usage.
+Given a gradient accumulation step of $n$, the model performs $n$ forward and backward passes while accumulating gradients before updating parameters.
+Since no parameter updates occur during this period, the gathered parameters can remain unsharded across multiple forward and backward passes.
+After a backward pass, the optimizer states are reloaded to GPU memory to update parameters.
 We run profiling after applying proactive prefetching to measure the peak memory usage.
 Then, we select as many parameters to keep unsharded as possible, ensuring that the total buffer size does not exceed the memory limit.
-The selection considers both the buffer size of each all-gather operator $o$, denoted as $\text{B}(o)$, and its communication time $T_{c}(\text{B}(o))$.
-Parameters are prioritized based on the ratio $\frac{T_{c}(\text{B}(o))}{\text{B}(o)}$, where a higher value indicates a greater reduction in communication time relative to memory cost.
-Following the trend shown in Fig.
-5(d), smaller parameters are generally selected first.
+The selection considers both the buffer size of each all-gather operator $o$, denoted as $\text{B}_{\text{ag}}(o)$, and its communication time $T_{c}(\text{B}_{\text{ag}}(o))$.
+Parameters are prioritized based on the ratio $\frac{T_{c}(\text{B}_{\text{ag}}(o))}{\text{B}_{\text{ag}}(o)}$, where a higher value indicates a greater reduction in communication time relative to memory cost.
+Since communication is less efficient for smaller messages, parameters with smaller buffer sizes are generally selected first.
 4.4. Adaptive offloading
-Although optimizer states such as momentum and variance used in Adam optimizer (Kingma and Ba, 2014) are only needed for parameter updates and are not required during the forward and backward passes, they occupy a significant amount of memory throughout training.
-Existing frameworks, such as DeepSpeed, provide functionality to offload these states to CPU memory.
-In fully sharded approaches like ZeRO-3, both model parameters and optimizer states are partitioned across GPUs.
-This design reduces the memory footprint on each device and enables parallel data transfers between GPUs and CPU memory.
-As a result, optimizer state offloading is often used together with fully sharded training.
+Optimizer states such as momentum and variance used in the Adam optimizer (Kingma and Ba, 2014) require a lot of GPU memory, but they are only needed for parameter updates and not during the forward or backward pass.
+Therefore, offloading them to CPU memory can significantly reduce peak memory usage.
+At the end of each backward pass, the optimizer states must be loaded back into GPU memory for parameter updates.
+As implemented in ZeRO-3, the fully sharded approach is often combined with offloading, as both techniques help reduce memory requirements when training large-scale models.
 Algorithm 2 Adaptive offloading (forward)
-0:  Initial schedule $S_{0}$, Memory limit $M$, optimizer state fragments $OS$
+0: Initial schedule $S_{0}$, Memory limit $M$, optimizer state fragments $OS=[os_{1},os_{2},\ldots]$
 $S\leftarrow[]$ // List of operators in output schedule
@@
 return $S$
-However, offloading and reloading optimizer states can typically significant overhead.
+However, data transfers for offloading and reloading optimizer states can incur significant overhead.
 To address this challenge, we propose an approach called adaptive offloading, which improves upon existing offloading mechanisms in two ways.
@@
 If the memory usage exceeds the limit, the algorithm adds operators to synchronize the copy and free memory for the fragments.
+As memory usage typically increases during the forward pass, this strategy enables efficient overlap of computation and offloading.
 We also iterate over the operators in the backward pass schedule, checking the available memory.
-If sufficient memory is available from that operator to the end, we schedule a reload operation to transfer the optimizer states back to the GPU.
+If sufficient memory is available from that operator to the end, we schedule an operator to initiate asynchronous transfer the optimizer states back to the GPU.
 This approach enables overlapping the backward computation with the data transfer from host memory to GPU memory while adhering to memory constraints.
-Since memory usage often increases during forward passes and decreases during backward passes, this strategy allows for efficient overlap of computation and offloading/reloading operations.
+In contrast to the forward pass, memory usage decreases during the backward pass, making it easier to hide reloading overhead.
+4.5. Composability
+DeepCompile organizes optimizations as modular compiler passes, enabling users to compose them flexibly based on their training goals and constraints.
+However, the effectiveness of these passes depends not only on their individual design but also on their application order and compatibility.
+The optimization passes for proactive prefetching, selective unsharding, and adaptive offloading are designed to be applied after the pass that enables the fully sharded approach, which rewrites the computation graph to insert all-gather and release operations for sharded parameters.
+However, the order and combination of these passes can significantly affect overall performance.
+A beneficial application order is to first apply proactive prefetching, followed by selective unsharding.
+This ordering ensures that selective unsharding can make use of the remaining memory after prefetching has opportunistically advanced communication while staying within memory constraints.
+In contrast, applying selective unsharding before proactive prefetching can reduce the effectiveness of prefetching.
+Since unsharding attempts to utilize as much available memory as possible to retain parameters in memory, little capacity remains for prefetch buffers.
+As a result, the benefits of prefetching are largely lost.
+Adaptive offloading is designed for scenarios where the model does not fit in GPU memory even after sharding.
+In practice, when adaptive offloading is applied, the memory available for proactive prefetching and selective unsharding is already limited, and the additional benefits from these passes tend to diminish.
 5. Evaluation
 We conducted experiments to evaluate the performance of DeepCompile across multiple dimensions, including computational efficiency, memory utilization, and correctness.
+Figure 7. Throughputs resulting from Llama-3 70B and Mixtral 8x7B models
 5.1. Experimental settings
-The experiments were performed on two or four servers, each equipped with eight NVIDIA H100 GPUs (80GB memory per GPU) and 1TB of system memory.
-The GPUs within each node were interconnected via NVLink, and each server included eight network interfaces (HCAs) connected via InfiniBand.
-The nodes were interconnected using InfiniBand with an observed bandwidth of 340GB/s for both allgather and reduce-scatter operations.
-Each server was equipped with two Intel(R) Xeon(R) Platinum 8462Y+ processors.
-In our experiments, we used the Llama-3 70B and Mixtral 8x7B models (47 billion parameters)(Jiang et al., 2024) as our target workloads.
-These models are representative examples of dense and Mixture-of-Experts (MoE) architectures(Shazeer et al., 2017; Lepikhin et al., 2021; Fedus et al., 2022), respectively.
-Both were run using bfloat16 with mixed precision (Micikevicius et al., 2017), while maintaining FP32 copies of parameters, gradients, and Adam optimizer states.
-We also employed activation checkpointing, integrated into the model implementation at the Transformer layer level, where each layer is recomputed as a block.
+The experiments were performed on two or four servers, each with eight NVIDIA H100 GPUs (80GB per GPU), 1TB of system memory, and two Intel(R) Xeon(R) Platinum 8462Y+ CPUs.
+GPUs within each server were connected via NVLink, and servers were interconnected using InfiniBand (340GB/s observed bandwidth for allgather and reduce-scatter).
+Each server also included eight InfiniBand-connected HCAs.
+We used Llama-3 70B and Mixtral 8x7B (47B) as representative dense and Mixture-of-Experts (MoE) models (Jiang et al., 2024; Shazeer et al., 2017; Lepikhin et al., 2021; Fedus et al., 2022).
+Both ran in bfloat16 with mixed precision (Micikevicius et al., 2017), keeping FP32 copies of parameters, gradients, and Adam states.
+We applied activation checkpointing (Chen et al., 2016) at the Transformer layer level, recomputing each layer as a block.
 We used PyTorch v2.6.0 and models from the HuggingFace Model Hub without any built-in parallelization support.
 The experiments were conducted with Python 3.10 and CUDA 12.4.
-DeepSpeed v1.6.4 and PyTorch FSDP were used as the baseline frameworks, and DeepCompile was built on top of DeepSpeed.
+ZeRO-3 (DeepSpeed v1.6.4) and PyTorch FSDP were used as the baseline frameworks, and DeepCompile was built on top of DeepSpeed.
 The optimization passes and profilers in DeepCompile were implemented in Python (approximately 2.6K lines), while the custom operators injected into the computational graph were implemented in C++ (approximately 900 lines).
-While some distributed training strategies support partial sharding—for example, sharding only optimizer states, or both gradients and optimizer states while keeping parameters unsharded—these configurations are not sufficient for our target models.
-The total parameter sizes of Llama-3 70B and Mixtral 8x7B exceed the memory capacity of a single GPU (140 GB and 94 GB, respectively), requiring all parameters, gradients, and optimizer states to be fully sharded across devices.
-For this reason, we use the fully sharded configurations of both DeepSpeed ZeRO-3 and PyTorch FSDP in our evaluation.
+Partial sharding strategies (e.g., sharding only optimizer states) are insufficient for our target models.
+Llama-3 70B and Mixtral 8x7B exceed single-GPU memory (140 GB and 94 GB), requiring all parameters, gradients, and optimizer states to be fully sharded.
+Thus, we use the fully sharded configurations of ZeRO-3 and FSDP in our evaluation.
 Other large-scale training frameworks, such as Megatron-LM, are not included in our comparison.
@@
 In contrast, we focus on model-agnostic approaches that can be applied to standard implementations.
-Similarly, frameworks that automate parallelization, such as Alpa (Mo et al., 2022), FlexFlow (Jia et al., 2018), and Unity (Unger et al., 2022), are not included, as they remain experimental and do not currently support the models used in this evaluation.
+Similarly, frameworks that automate parallelization, such as Alpa (Mo et al., 2022), FlexFlow (Jia et al., 2019), nnScalar (Lin et al., 2024), and Unity (Unger et al., 2022), are not included, as they do not currently support the models used in this evaluation.
 5.2. Efficiency
 Figure 7 presents the throughputs observed in our experiments.
-We used Llama-3-70B with 32 GPUs and Mixtral-8x7B with 16 GPUs.
-We evaluated DeepSpeed ZeRO-3 both without and with the PyTorch compiler enabled, labeled as ZeRO3 and ZeRO3 (C), respectively.
-Similarly, we include results for PyTorch FSDP, labeled as FSDP and FSDP (C) depending on whether the PyTorch compiler is enabled.
+We used Llama-3-70B and Mixtral-8x7B with 32 GPUs.
+All throughput numbers reflect end-to-end time per training iteration, including the forward and backward passes as well as parameter updates.
+We evaluated ZeRO-3 both without and with the PyTorch compiler enabled, labeled as ZeRO3 and ZeRO3 (C), respectively.
+Similarly, we include results for FSDP, labeled as FSDP and FSDP (C) depending on whether the PyTorch compiler is enabled.
 For DeepCompile, we tested three configurations: enabling only proactive prefetching, enabling only selective unsharding, and enabling both in that order.
@@
 Since the models fit within the available GPU memory at these scales, adaptive offloading was not applied in these experiments.
-We conducted this experiment using different batch sizes and sequence lengths. In addition, we evaluated training performance with gradient accumulation.
-Figure 7. Throughputs resulting from Llama-3 70B and Mixtral 8x7B models
-DeepCompile consistently outperformed all baseline methods, with the most significant gains observed when both proactive prefetching and selective unsharding were enabled (DeepCompile (P+S)).
-Under the setting with a gradient accumulation step of 1, the smallest improvement over ZeRO-3 was 1.11×, observed in the configuration with sequence length 512 and batch size 1.
-In this setting, computation is significantly faster relative to communication, leaving little opportunity for our optimizations to hide communication overhead through overlap.
-As a result, the benefit of prefetching and unsharding becomes limited.
-For all other conditions, DeepCompile achieved approximately 1.14× to 1.18× speedup over ZeRO-3, and 1.16× to 1.26× over FSDP in the Llama-3-70B experiments.
-For Mixtral-8x7B, the speedups over ZeRO-3 ranged from 1.07× to 1.35×, while those over FSDP ranged from 1.06× to 1.14×.
+We conducted experiments with varying batch sizes, sequence lengths, and gradient accumulation steps.
+The per-GPU batch size (number of sequences) was set to 1, 2, or 4, and the sequence length to 512, 1024, or 2048.
+We also varied the gradient accumulation steps from 1 to 16.
+The effective batch size, measured in tokens, is given by: $\text{\#Sequences per GPU}\times\text{Sequence length}\times\text{Accumulation steps}\times\text{\#GPUs}$.
+Under these settings, the total number of tokens per step ranged from 8K to 4M, covering a broad range representative of common configurations used for training models at this scale.
+For the system parameters defined in Table 1, we set $M$ to 90% of the available GPU memory to maintain a safety margin.
+$M_{\text{prefetch}}$ was set to 2 GB.
+Increasing this value allows DeepCompile to schedule more all-gather communications earlier, but we did not observe any clear performance benefit from setting it higher.
+The parameter $\alpha$ controls the aggressiveness of all-gather fusion.
+While larger values encourage more aggressive fusion, we found that increasing $\alpha$ beyond 1.5 had little effect on performance.
+Based on experiments with values ranging from 1.0 to 2.0, we fixed $\alpha$ to 1.5.
+We used these values throughout all of the experiments.
+DeepCompile consistently outperformed all baselines, with the largest gains observed when both proactive prefetching and selective unsharding were enabled (DeepCompile (P+S)).
+With gradient accumulation set to 1, the smallest improvement over ZeRO-3 for Llama-3-70B (Fig 7(a),(i)(ii)) was 1.11×, observed at sequence length 512 and batch size 1.
+In this configuration, computation dominates communication, leaving limited room for our optimizations to hide communication latency via overlap.
+Consequently, the benefits of prefetching and unsharding are reduced.
+Under all other conditions, DeepCompile achieved a 1.14×–1.18× speedup over ZeRO-3 and 1.16×–1.26× over FSDP.
+For Mixtral-8x7B, the speedups over ZeRO-3 ranged from 1.07× to 1.35×, while those over FSDP ranged from 1.06× to 1.14× (Figure 7(b),(i)(ii)).
 Mixtral-8x7B is an MoE model, where only a subset of parameters is activated for each input token.
@@
 Figure 7(a)(iii) and (b)(iii) show throughput results when varying the gradient accumulation steps from 1 to 16.
+All numbers reflect the end-to-end time per training iteration, including all gradient accumulation steps.
 As discussed in Section 4.3, this setting significantly benefits from selective unsharding.
@@
 5.4. Adaptive Offloading
-The Llama-3-70B model requires 32 or more GPUs, each with 80GB of memory, to fit entirely into GPU memory under our current settings.
-However, by offloading optimizer states, we are able to fit the model onto just 16 GPUs.
-To evaluate the benefits of adaptive offloading of optimizer states, as described in Section 4.4, we compared the throughput achieved using DeepSpeed’s offloading feature with that of our adaptive offloading approach.
-Figure 9 shows the iteration times for each method.
-ZeRO3 (Offload optimizer) and ZeRO3 (Offload optimizer) + Compile represent the results of using DeepSpeed ZeRO-3 with and without the PyTorch compiler, respectively.
-Notably, DeepSpeed not only offloads optimizer states to host memory but also performs parameter updates on the CPU.
+The Llama-3-70B model requires at least 32 GPUs (each with 80GB memory) to fit entirely into GPU memory under our settings.
+However, by offloading optimizer states, we are able to run the model on just 16 GPUs.
+To evaluate the benefits of adaptive optimizer offloading (Section 4.4), we compared the throughput achieved by DeepSpeed’s offloading mechanism combined with ZeRO-3 against our adaptive offloading approach.
+Figure 9 presents the end-to-end iteration time for each method.
+ZeRO3 (Offload optimizer) and ZeRO3 (Offload optimizer) + Compile represent the results of using ZeRO-3 with and without the PyTorch compiler, respectively.
+Notably, ZeRO-3 not only offloads optimizer states to host memory but also performs parameter updates on the CPU.
 While this approach saves more memory than DeepCompile ’s offloading, it significantly slows down parameter updates.
-Since this behavior makes it difficult to compare DeepSpeed’s offloading directly with our method—which keeps parameter updates on the GPU—we also included results for DeepCompile (Offload optimizer all + sync), which offloads all optimizer states at the beginning of the forward pass and reloads them synchronously at the end.
+Since this behavior makes it difficult to compare ZeRO-3’s offloading directly with our method, which keeps parameter updates on the GPU, we also included results for DeepCompile (Offload optimizer all + sync), which offloads all optimizer states at the beginning of the forward pass and reloads them synchronously at the end.
 Our adaptive offloading method is shown as DeepCompile (Offload selective + async), which minimizes the amount of data offloaded and overlaps offloading and reloading asynchronously with computation.
@@
 In our experiments, we first executed the pass that inserts allgather and release operations.
-After a short warmup period (five training iterations), we applied additional passes including proactive prefetching and selective unsharding.
+After a short warmup period (five training iterations in out experiments), we applied additional passes including proactive prefetching and selective unsharding.
 These passes involve graph analysis and profiling, resulting in several minutes of overhead for our target models.
@@
 Mixtral 8x7B 437.0 s 416.5 s 407.8 s
-The variation in compile time across optimization configurations for the same model is less than 10%, indicating that the additional cost of enabling proactive prefetching or selective unsharding is relatively minor compared to the base analysis and scheduling pass.
+The variation in compile time across optimization configurations for the same model is less than 10%, indicating that the additional cost of enabling proactive prefetching or selective unsharding is relatively minor compared to the base analysis and the pass to enable the fully-shareded approach.
 On the other hand, compile times vary more significantly across models, as seen in the longer durations for Mixtral-8x7B compared to Llama-3-70B.
@@
 5.6. Correctness
-To verify the correctness of the optimization passes in DeepCompile, we compared the resulting loss values with those of DeepSpeed ZeRO-3.
+To verify the correctness of the optimization passes in DeepCompile, we compared the resulting loss values with those of ZeRO-3.
 We initialized Llama-3 70B model with random weights but the same random seed in both settings.
@@
 Figure 10. Verification of loss values
+Table 3. Comparison of DeepCompile with existing distributed training systems
+Feature ZeRO-3 / FSDP (Rajbhandari et al., 2020; Team, 2023) SimpleFSDP (Zhang et al., 2024) Alpa / Unity (Mo et al., 2022; Unger et al., 2022) DeepCompile (ours)
+No code modification
+Compiler-level optimization
+Profiling guided optimization
+Communication overlap
+Composable optimizations
+Memory-aware scheduling
 6. Related Work
-6.1. Distributed Training Framework
-As discussed in Section 2, various distributed training strategies have been proposed to scale deep learning models beyond the capacity of a single GPU.
-For instance, tensor parallelism, as introduced in Megatron-LM (Shoeybi et al., 2019), splits large matrix operations across devices, while pipeline parallelism, as implemented in GPipe (Huang et al., 2019), assigns consecutive layers to different GPUs.
-While effective, both approaches require the user to manually modify the model to specify how tensors or layers are partitioned.
-This makes it challenging to adapt these strategies to diverse or rapidly evolving model architectures.
-In contrast, our work focuses on the fully sharded approach, as implemented in ZeRO (Rajbhandari et al., 2020) and FSDP (Team, 2023), which requires no model code changes and is applicable to a wide range of architectures.
-DeepCompile builds on this fully sharded approach and introduces compiler-level optimizations to improve its efficiency.
-Note that DeepCompile is not inherently limited to the fully sharded approach.
-Its design is compatible with other parallelization strategies.
-For example, RaNNC (Tanaka et al., 2021) automatically partitions the computation graph to enable pipeline parallelism, and Alpa (Mo et al., 2022) selects distributed execution rules based on the partitioned state of tensors.
-Similarly, DeepCompile could support automated parallelization by replacing computation operators in the graph and integrating this with other optimization passes within a unified framework.
-6.2. Optimizing Parallelization Strategies
-To improve training efficiency for large-scale models, various strategies such as data, tensor, and pipeline parallelism have been studied.
-Recent work has focused on automatically discovering optimal combinations of these strategies, including Alpa (Mo et al., 2022), FlexFlow (Jia et al., 2018), Unity (Unger et al., 2022), and Galvatron (Miao et al., 2022).
-Dynamic switching techniques have also been proposed, such as HotSpa (Ge et al., 2024), which aims to adaptively reconfigure parallelization strategies during training.
-However, these approaches generally assume that the model is already implemented in a way that supports the desired forms of parallelism.
-In practice, tensor and pipeline parallelism often require model-specific partitioning logic and integration with low-level APIs.
-Supporting multiple strategies and enabling dynamic switching can involve considerable engineering effort, particularly for complex or evolving model architectures.
-As a result, despite their promise, such automated parallelization systems have seen limited adoption in current practice.
-6.3. Computation-Communication Overlapping
-As discussed earlier, widely used frameworks such as ZeRO-3 and FSDP incorporate prefetching techniques to overlap communication with computation and reduce latency from all-gather operations.
-Prefetching has become a standard optimization for improving efficiency in fully sharded training.
-More recently, techniques for overlapping communication at a finer granularity, such as intra-layer parallelism, have also been proposed (Wang et al., 2022, 2024).
-These approaches aim to eliminate or hide communication costs by restructuring computation within layers.
-In contrast, our work focuses on profiling-guided, graph-level optimization of communication overlap.
-Like DeepCompile, SimpleFSDP (Zhang et al., 2024) targets compiler-level optimization of the fully sharded approach.
-It performs memory-aware fusion of all-gather operations to reduce communication overhead, but does not take into account dynamic changes in memory usage when deciding how early each all-gather can be issued without exceeding memory constraints.
-Moreover, SimpleFSDP focuses solely on scheduling communication for FSDP.
-In contrast, DeepCompile is designed to support a broader range of optimizations and enables techniques such as selective unsharding and adaptive offloading.
+Distributed training has become a central topic in deep learning system design, leading to a range of frameworks that support model parallelism, memory optimization, and scaling strategies.
+We summarize key systems and contrast their capabilities with our approach in Table 3.
+Framework-level sharding Fully sharded training frameworks such as ZeRO-3 (Rajbhandari et al., 2020) and PyTorch FSDP (Team, 2023) partition model parameters and optimizer states across devices and apply runtime optimizations like prefetching and unsharding.
+These optimizations are typically implemented using hooks inserted at the framework level and rely on heuristics that do not leverage global analysis or profiling feedback.
+As a result, they offer limited control over the scheduling of communication and memory operations, and cannot coordinate multiple interacting optimizations in a globally informed manner.
+Compiler-based prefetching SimpleFSDP (Zhang et al., 2024) uses a compiler to improve communication scheduling for fully sharded training.
+It analyzes operator-level memory usage to decide whether to fuse multiple all-gather operations.
+However, it focuses solely on prefetching and does not support coordination across multiple optimizations or feedback loops between passes.
+In addition, while it considers memory usage for fusion decisions, it does not take memory dynamics into account when determining the placement of all-gather operations.
+DeepCompile generalizes this idea by introducing a pass-based architecture in which each optimization pass can access updated profiling results and respond to memory dynamics.
+Automatic parallelization planning FlexFlow (Jia et al., 2019) proposed selecting optimal combinations of data, tensor, and pipeline parallelism based on cost models.
+Building on this idea, recent systems such as Alpa (Mo et al., 2022), nnScaler (Lin et al., 2024), and Unity (Unger et al., 2022) integrate parallelism planning with compiler-based optimization techniques.
+These systems focus on cost-model-driven planning and static graph analysis to combine parallelism strategies and GPU-local optimizations.
+However, they do not account for runtime memory usage trends or support techniques such as communication–computation overlap and memory-aware offloading.
+DeepCompile complements these approaches by emphasizing profiling-guided, memory-centric optimization using modular passes that adapt to runtime behavior.
+Although our current focus is on the fully sharded approach, DeepCompile can also express other parallelism strategies.
+For example, tensor parallelism can be realized by identifying Transformer blocks and injecting all-reduce operations; pipeline parallelism can be enabled by partitioning the graph into stages with balanced compute and memory load; and sequence parallelism (Jacobs et al., 2023; Liu et al., 2023) can be supported through all-to-all or peer-to-peer communication.
+A key distinction is that DeepCompile’s modular pass framework enables coordinated composition of such parallelization strategies with memory-aware techniques like offloading.
+This interaction has received little attention in prior work and represents a promising direction for future research.
 7. Conclusion
-We introduced DeepCompile, a compiler-based framework for optimizing distributed training of large-scale deep learning models.
-DeepCompile starts from a standard single-GPU computation graph and incrementally inserts distributed operators such as all-gather and release, enabling graph-level analysis and holistic optimization of distributed training.
-Unlike existing frameworks like ZeRO-3 and FSDP, which rely on heuristics and fixed module boundaries, DeepCompile leverages runtime profiling to flexibly reorder and coordinate operations based on actual memory usage and data dependencies.
-Built on this foundation, DeepCompile implements a fully sharded training approach along with three key optimizations: proactive prefetching to improve communication–computation overlap, selective unsharding to reduce redundant communication, and adaptive offloading to dynamically manage memory under tight resource constraints.
-These optimization passes are applied in sequence, informed by profiling feedback, allowing DeepCompile to coordinate them effectively in a unified compilation framework.
-Our experiments demonstrate that DeepCompile consistently outperforms ZeRO-3 and FSDP baselines across a range of configurations.
-It achieves up to 1.28× and 1.54× higher throughput for Llama-3 70B and Mixtral 8×7B models, respectively.
-Moreover, in scenarios with limited GPU memory, DeepCompile’s adaptive offloading achieves up to a 7.01× throughput improvement by automatically overlapping communication and reducing memory pressure.
+We presented DeepCompile, a compiler-based optimization framework for distributed deep training.
+By transforming user-defined models into computation graphs and applying a series of profiling-guided optimization passes, DeepCompile enables fine-grained memory-aware scheduling and communication planning across passes.
+This approach contrasts with existing systems like ZeRO-3 and FSDP, which lack flexibility in communication timing and memory management.
+Built on a fully sharded foundation, DeepCompile incorporates three key optimization passes: proactive prefetching to improve communication–computation overlap, selective unsharding to reduce communication, and adaptive offloading to manage memory when resources are limited.
+Each pass utilizes updated profiling feedback, allowing the system to compose optimizations based on their combined effects.
+Our evaluation demonstrates that DeepCompile consistently outperforms existing baselines.
+It achieves up to 1.28× and 1.54× throughput improvements on Llama 3 70B and Mixtral 8×7B models, respectively.
+In memory-constrained settings, adaptive offloading enables up to 7.01× throughput gains by overlapping data transfers with computation.
 DeepCompile represents a step forward in compiler-driven distributed training.
-Its graph transformation capabilities and runtime-aware optimization passes offer a scalable and automated path toward efficient parallelization.
-In future work, we aim to extend DeepCompile to support broader classes of optimizations, such as fully automated parallelization planning and enhanced memory scheduling.
+Its graph transformation capabilities and runtime-aware optimization passes provide a scalable and automated path to efficient parallelization.
+As discussed in Section 6, future work will extend DeepCompile to support a broader class of optimizations, including communication and memory scheduling.